### 16 Mar 2026 v1

# On the first-order genus of wreath products and their central extensions

Abstract
We prove that groups of the form $\mathbb{Z}^{m}\,{\rm wr}\,\mathbb{Z}^{n}$, where $m,n\in\mathbb{N}$, are regularly bi-interpretable with $\mathbb{Z}$ and therefore are first-order rigid: every finitely generated group elementarily equivalent to $\mathbb{Z}^{m}\,{\rm wr}\,\mathbb{Z}^{n}$ is isomorphic to $\mathbb{Z}^{m}\,{\rm wr}\,\mathbb{Z}^{n}$. On the other hand, we show that $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$ admits $2^{\aleph_{0}}$ elementarily equivalent, pairwise non-isomorphic central extensions with finite kernel.
Contents
1 Introduction
Classification of groups with respect to elementary equivalence is an area of algebra and mathematical logic that goes back to the pioneering works of Löwenheim, Skolem, and Tarski in the 1920s. The Löwenheim-Skolem theorem tells us that for a given infinite group $G$ and any infinite cardinal $\lambda$, there is a group $H$ of cardinality $\lambda$ that is elementarily equivalent to $G$ (symbolically, $G\equiv H$). Hence, the first-order theory $Th(G)$ of $G$ – that is, the set of all sentences in the language of group theory $\{\cdot,{}^{-1},1\}$ that are true in $G$ – does not uniquely characterize the group $G$ up to isomorphism. This observation gives rise to the first-order classification problem, originally formulated by Tarski: for a given group $G$, characterize all groups elementarily equivalent to $G$.
This problem turned out to be difficult, so a more tractable question was posed: how many groups $H$ of infinite cardinality $\lambda$ satisfy $H\equiv G$?
The extremal instance of the Löwenheim-Skolem theorem occurs when there exists, up to isomorphism, a unique group $H$ of cardinality $\lambda$ such that $G\equiv H$. Morley termed such theories $Th(G)$ $\lambda$-categorical and proved that if $Th(G)$ is categorical in some uncountable cardinal, then it is categorical in all uncountable cardinals (and he did so for arbitrary algebraic structures, not just groups).
Unfortunately, there are not many groups that are uncountably or countably categorical. Furthermore, if the theory $Th(G)$ is unstable, then it follows from Shelah’s classification theory that the number of non-isomorphic models of a given uncountable cardinal $\kappa$ is maximal, which is $2^{\kappa}$.
[…]
Therefore, when viewed through the lens of cardinalities, the standard model-theoretic approach to models of complete theories is, in general, not very informative. A recent promising development in the first-order classification problem arises from within group theory itself. From an algebraic viewpoint, it is natural to consider all finitely generated (rather than merely countable) groups $H$ such that $G\equiv H$. To this end, for a finitely generated group $G$ we define the first-order genus $\mathcal{FO}(G)$ as the set of isomorphism classes $[H]$ of all finitely generated groups $H$ with $G\equiv H$.
In [3] Avni, Lubotzky, and Mieri termed a finitely generated group $G$ first-order rigid if for any finitely generated group $H$, if $G\equiv H$ then $G\simeq H$. In other words, $G$ is first-order rigid if and only if $|\mathcal{FO}(G)|=1$. It turned out that there are many first-order rigid groups, despite the fact that there are only a few countably categorical ones, and there are no infinite, finitely generated, countably categorical groups at all [6]. Curiously, among first-order rigid groups, there are some that can be completely characterized by a single sentence in the language of group theory. Following Nies [30, 31], a finitely generated group $G$ is termed quasi-finitely axiomatizable (QFA) if there exists a sentence $\sigma$ in the language of group theory such that, whenever $\sigma$ holds in a finitely generated group $H$, we have $G\simeq H$.
Note that finite groups $G$ are not only QFA but also finitely axiomatizable; that is, any group elementarily equivalent to $G$ is in fact isomorphic to $G$. Infinite finitely generated abelian groups are first-order rigid but not QFA. The former is straightforward, while the latter follows from Baur–Monk quantifier elimination for abelian groups (see, for example, [38]).
If $G$ is a finitely generated nilpotent group, then $\mathcal{FO}(G)$ is finite. To see this note that due to Pickel’s result from [36] there are only finitely many up to isomorphism finitely generated nilpotent groups which have the same (up to isomorphism) finite quotients as the group $G$. This set of isomorphism classes is called the profinite genus of $G$. Now, for a finitely generated group $H$, if $G\equiv H$, then $H$ is nilpotent and the quotients $G/G^{k}G_{i}$ and $H/H^{k}H_{i}$, where $G_{i},H_{i}$ are $i$-th terms of the lower central series, and $G^{k}$ and $H^{k}$ are the subgroups generated by $k$-th powers of elements in $G$ and $H$, are elementarily equivalent for each $k$ and $i$ (because verbal subgroups in finitely generated nilpotent groups have finite width, hence definable in these groups by the same formulas) and finite, hence isomorphic.
Furthermore, most finitely generated nilpotent groups $G$ are QFA. Indeed, Oger and Sabbagh showed in [33] that a finitely generated nilpotent group $G$ is QFA if and only if the center of $G$ is contained in the isolator of the commutant of $G$ and due to [14] a random finitely generated nilpotent group satisfies this property. A similar line of reasoning works for polycyclic groups $G$: Sabbagh and Wilson showed in [42] that if $H$ is a finitely generated group and $G\equiv H$ then $H$ is polycyclic; Oger established in [32] that if $G$ and $H$ are polycyclic groups and $G\equiv H$ then $G$ and $H$ have the same finite quotients; and Grunewald, Pickel, and Segal proved in [12] that there are only finitely many up to isomorphism polycyclic groups $H$ which have the same (up to isomorphism) finite quotients as the group $G$. Hence, for a polycyclic group $G$, the set $\mathcal{FO}(G)$ is finite. Similarly to nilpotent groups, Lasserre described in [23] all QFA polycyclic groups.
However, this line of reasoning fails for finitely generated metabelian groups, since Pickel showed in [37] that there exist infinitely many pairwise non-isomorphic, finitely presented metabelian groups sharing the same finite quotients. For finitely generated metabelian groups $G$ that are not polycyclic, little is known about the conditions under which they are elementarily equivalent, about their first-order genus, or about related model-theoretic properties.
Khelif showed in [19] that the metabelian Baumslag-Solitar groups $BS(1,n)$ for $n>1$, as well as the free metabelian group $M_{2}$ of rank 2, are QFA. In [20], Kharlampovich and Myasnikov proved that all non-abelian free metabelian groups $M_{n}$ of finite rank $n$ are QFA, and they provided a complete description of all groups (not necessarily finitely generated) that are elementarily equivalent to $M_{n}$. Subsequently, in [9], Danyarova and Myasnikov described all groups $H$ such that $H\equiv BS(1,n)$. All these results for metabelian non-polycyclic groups $G$ are obtained via a novel method based on bi-interpretations of $G$ with standard arithmetic $\mathbb{N}$, or equivalently, with the ring of integers $\mathbb{Z}$.
A more general approach to the QFA property in groups, grounded in the richness of their algebraic structure, is introduced in [22]. An algebraic structure is called rich if its first-order theory has the same expressive power as its weak second-order theory. It was shown in [22] that rich groups with an arithmetic multiplication table are QFA, and also that groups regularly bi-interpretable with $\mathbb{Z}$ are rich and have an arithmetic multiplication table, hence QFA. There are many rings and groups that are bi-interpretable with $\mathbb{Z}$. Examples include finitely generated infinite integral domains [2]; finitely generated fields (with a few exceptions) [10]; Chevalley groups (with some exceptions) over such rings [7]; Thompson’s groups $F$ and $T$ [24]; and centerless, non-uniform, irreducible higher-rank arithmetic lattices in characteristic zero [4]. We refer to [22, 8] for details and further examples. On the other hand we mention here that free non-abelian groups $F$ are not QFA; $|\mathcal{FO}(F)|=\kappa_{0}$; and all finitely generated groups $H$ with $F\equiv H$ have been completely described [21, 44].
In this paper we prove two main results. The first is the following theorem, which answers Problem 4.3 from [22] in the affirmative.
Theorem 1.1.
Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is regularly injectively bi-interpretable with $\mathbb{Z}$.
The proof of Theorem 1.1 is divided into four parts: interpretability of $G$ in $\mathbb{Z}$} in Section 3.1, interpretability of $\mathbb{Z}$} in $G$ in Sectiom 3.2, bi-interpretability of $G$ and $\mathbb{Z}$} with parameters, and finally, regular bi-interpretability of $G$ and $\mathbb{Z}$} in Sections 4, 5.
Corollary 1.2.
Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is QFA.
Theorem 1.1 and results from [8], Section 5.8 imply the following.
Corollary 1.3.
[…]
Our second result shows (see below) that there exists a finite central extension $G$ of $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$ such that $|\mathcal{FO}(G)|=2^{\aleph_{0}}$. To state the second theorem, we need some notation and terminology.
Recall that the space of finitely generated marked groups $\mathcal{G}$ can be informally defined as the set of all pairs $(G,A)$, where $G$ is a group and $A$ is an ordered finite generating set of $G$, endowed with the topology induced by local convergence of Cayley graphs. Given a finitely generated group $G$, we denote by $G]\subseteq\mathcal{G}$ its isomorphism class; that is,
$G]=\{(H,B)\in\mathcal{G}\mid H\cong G\}.$
The following definition is inspired by connections between the topological properties of isomorphism classes in $\mathcal{G}$ and model theory (see [34]).
Definition 1.4.
A finitely generated group $G$ is said to be condensed if $[G]$ has no isolated points.
We prove the following theorem.
Theorem 1.5.
There exists a condensed group $G$ that splits as a central extension of the form
$1\longrightarrow\mathbb{Z}_{2}\longrightarrow G\longrightarrow\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}\longrightarrow 1.$ (1)
There are three immediate consequences of this result. The first one concerns the complexity of the isomorphism relation. Recall that an equivalence relation $E$ on a topological space $X$ is called smooth if there is a Polish space $P$ and a Borel map $\beta\colon X\to P$ such that for any $x,y\in X$, we have $xEy$ if and only if $\beta(x)=\beta(y)$.
In [46], Williams proved that the isomorphism relation on the space of $3$-step solvable, finitely generated, marked groups is non-smooth. We strengthen this result as follows.
Corollary 1.6.
The isomorphism relation on the space of center-by-metabelian, finitely generated, marked groups is not smooth.
Informally, our result means that finitely generated center-by-metabelian groups cannot be “explicitly classified” up to isomorphism using invariants from a Polish space. Note that Corollary 1.6 is optimal in a certain sense. Indeed, the isomorphism relation on the space of finitely generated marked metabelian groups is obviously smooth, as the latter space is countable.
Another corollary answers the second part of Question 7.2 in [34].
Corollary 1.7.
The property of being condensed is not geometric, i.e., not stable under quasi-isometry of finitely generated groups.
The last corollary sharply contrasts the fact that $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$ is QFA.
Corollary 1.8.
There exist $2^{\aleph_{0}}$ finitely generated, pairwise non-isomorphic, elementarily equivalent central extensions of the form (1).
2 Preliminaries
2.1 Interpretability and bi-interpretability
One can use the model-theoretic notion of interpretability and bi-interpretability to study structures elementarily equivalent to a given one. In this paper we are going to do this for the restricted wreath product ${\mathbb{Z}}^{n}\,{\rm wr}\,{\mathbb{Z}}^{m}$. We remind here some precise definitions and several known facts that may not be very familiar to algebraists.
Let $\mathbb{B}=\langle B;L\rangle$ be a structure. A subset $A\subseteq B^{n}$ is called definable in $\mathbb{B}$ if there is a formula $\phi(x_{1},\ldots,x_{n})$ (without parameters) in $L(\mathbb{B})$ such that
$A=\{(b_{1},\ldots,b_{n})\in B^{n}\mid\mathbb{B}\models\phi(b_{1},\ldots,b_{n})\}.$
In this case we denote $A$ by $\phi(B^{n})$ or $\phi({\mathbb{B}})$ and say that $\phi$ defines $A$ in $\mathbb{B}$. Similarly, an operation $f$ on the subset $A$ is definable in $\mathbb{B}$ if its graph is definable in $\mathbb{B}$. A constant $c$ is definable if the relation $x=c$ is definable. An $n$-ary predicate $P(x_{1},\ldots,x_{n})$ is definable in $\mathbb{B}$ if the set $\{(b_{1},\ldots,b_{n})\in\mathbb{B}^{n}|P(b_{1},\ldots,b_{n})\ {\rm is\ true}\}$ is definable in $\mathbb{B}$.
In the same vein an algebraic structure $\mathbb{A}=\langle A;f,\ldots,P,\ldots,c,\ldots\rangle$ is definable in $\mathbb{B}$ if there is a definable subset $A^{*}\subseteq B^{n}$ and operations $f^{*},\ldots,$ predicates $P^{*},\ldots,$ and constants $c^{*},\ldots,$ on $A^{*}$ all definable in $\mathbb{B}$ such that the structure $\mathbb{A}^{*}=\langle A^{*};f^{*},\ldots,P^{*},\ldots,c^{*},\ldots,\rangle$ is isomorphic to $\mathbb{A}$. (Note that constants $c,\ldots$ belong to the language of $\mathbb{A}$, they are not parameters.) For example, the center of any group $G$ (considered as a structure in the standard group theoretic language) is definable as a group in $G$.
In the notation above if $\sim$ is a definable equivalence relation on a definable subset $A\subseteq B^{n}$ then we say that the quotient set $A/\sim$ is interpretable in $\mathbb{B}$. Furthermore, an operation $f$ or a predicate $P$ on the quotient set $A/\sim$ is interpretable in $\mathbb{B}$ if the full preimage of its graph in $A$ is definable in $\mathbb{B}$. For example, if $N$ is a normal definable subgroup of a group $G$, then the equivalence relation $x\sim y$ on $G$ given by $xN=yN$ is definable in $G$, so the quotient set $G/N$ of all right cosets of $N$ is interpretable in $G$. It is easy to see that the multiplication induced from $G$ on $G/N$ is also interpretable in $G$. This shows that the quotient group $G/N$ is interpretable in $G$.
Definition 2.1.
An algebraic structure $\mathbb{A}=\langle A;f,\ldots,P,\ldots,c,\ldots\rangle$ is absolutely interpretable (or 0-interpretable) in a structure $\mathbb{B}$ if there is a subset $A^{*}\subseteq B^{n}$ definable in $\mathbb{B}$, an equivalence relation $\sim$ on $A^{*}$ definable in $\mathbb{B}$, operations $f^{*},\ldots,$ predicates $P^{*},\ldots,$ and constants $c^{*},\ldots,$ on the quotient set $A^{*}/{\sim}$ all interpretable in $\mathbb{B}$ such that the structure $\mathbb{A}^{*}=\langle A^{*}/{\sim};f^{*},\ldots,P^{*},\ldots,c^{*},\ldots,\rangle$ is isomorphic to $\mathbb{A}$.
Now we introduce some notation. An interpretation of $\mathbb{A}$} in $\mathbb{B}$} is described by the following set of formulas in the language $L({\mathbb{B}})$
$\Gamma=\{U_{\Gamma}(\bar{x}),E_{\Gamma}(\bar{x}_{1},\bar{x}_{2}),Q_{\Gamma}(\bar{x}_{1},\ldots,\bar{x}_{t_{Q}})\mid Q\in L({\mathbb{A}})\}$
(here $\bar{x}$ and $\bar{x}_{i}$ are $n$-tuples of variables) which interpret $\mathbb{A}$ in $\mathbb{B}$ (as in the definition 2.1 above). Namely, $U_{\Gamma}$ defines in $\mathbb{B}$ a subset $A_{\Gamma}=U_{\Gamma}(B^{n})\subseteq B^{n}$, $E_{\Gamma}$ defines in $\mathbb{B}$} an equivalence relation $\sim_{\Gamma}$ on $A_{\Gamma}$, and the formulas $Q_{\Gamma}$ define functions $f_{\Gamma}$, predicates $P_{\Gamma}$, and constants $c_{\Gamma}$ that interpret the corresponding symbols from $L(\mathbb{A})$ on the quotient set $A_{\Gamma}/\sim_{\Gamma}$ in such a way that the $L$-structure $\Gamma({\mathbb{B}})=\langle A_{\Gamma}/\sim_{\Gamma};f_{\Gamma},\ldots,P_{\Gamma},\ldots,c_{\Gamma},\ldots\rangle$ is isomorphic to $\mathbb{A}$}. Note, that we interpret a constant $c\in L({\mathbb{A}})$ in the structure $\Gamma({\mathbb{B}})$ by the $\sim_{\Gamma}$-equivalence class of some tuple $\bar{b}_{c}\in A_{\Gamma}$ defined in $\mathbb{B}$} by the formula $Q_{c}$. We write $\mathbb{A}\simeq\Gamma(\mathbb{B})$ if $\Gamma$ interprets $\mathbb{A}$} in $\mathbb{B}$} as described above and refer to $\Gamma$ as an interpretation code or just code. The number $n$ is called the dimension of $\Gamma$, denoted $n=dim\Gamma$. By $\mu_{\Gamma}$ we denote a a surjective map $A_{\Gamma}\to{\mathbb{A}}$ (here ${\mathbb{A}}=\langle A;L({\mathbb{A}})\rangle$) that gives rise to an isomorphism $\bar{\mu}_{\Gamma}:\Gamma({\mathbb{B}})\to\mathbb{A}$. We refer to this map $\mu_{\Gamma}$ as the the coordinate map of the interpretation $\Gamma$. Sometimes we cal the relation $\sim_{\Gamma}$ the kernel of the coordinate map $\mu_{\Gamma}$ and denote it by $\ker(\mu_{\Gamma})$. Finally, notation $\mu:{\mathbb{B}}\rightsquigarrow{\mathbb{A}}$ means that $\mathbb{A}$} is interpretable in $\mathbb{B}$} with the coordinate map $\mu$. We use this notation throughout the paper.
More generally, the formulas that interpret $\mathbb{A}$ in $\mathbb{B}$ may contain elements from $\mathbb{B}$ that are not in the language $L(\mathbb{B})$, i.e., some parameters, say $p_{1},\ldots,p_{k}\in B$. In this case we assume that all the formulas from the code $\Gamma$ have a tuple of extra variables $\bar{y}=(y_{1},\ldots,y_{k})$ for parameters in $\mathbb{B}$}:
$\Gamma=\{U_{\Gamma}(\bar{x},\bar{y}),E_{\Gamma}(\bar{x}_{1},\bar{x}_{2},\bar{y}),Q_{\Gamma}(\bar{x}_{1},\ldots,\bar{x}_{t_{Q}},\bar{y})\mid Q\in L(\mathbb{A})\}$ (2)
so that after the assignment $y_{1}\to p_{1},\ldots,y_{k}\to p_{k}$ the code interprets $\mathbb{A}$ in $\mathbb{B}$. In this event we write $\mathbb{A}\simeq\Gamma(\mathbb{B},\bar{p})$ (here $\bar{p}=(p_{1},\ldots,p_{k})$), and say that $\mathbb{A}$ is interpretable in $\mathbb{B}$ by the code $\Gamma$ with parameters $\bar{p}$. In the case when $\bar{p}=\emptyset$ one gets again the absolute interpretability.
We will say that a subset $D\subseteq A_{\Gamma}/\sim_{\Gamma}$ is definable in $\mathbb{B}$} if its full preimage in $A_{\Gamma}$ is definable in $\mathbb{B}$}. More generally, a subset $D\subseteq(A_{\Gamma}/\sim_{\Gamma})^{m}$ is definable in $\mathbb{B}$} if its full preimage in $A_{\Gamma}^{m}$ under the natural projection $A_{\Gamma}^{m}\to(A_{\Gamma}/\sim_{\Gamma})^{m}$ is definable in $\mathbb{B}$}.
Now we discuss a very strong version of mutual interpretability of two structures, so-called bi-interpretability.
Definition 2.2.
Two algebraic structures $\mathbb{A}$} and $\mathbb{B}$} are called bi-interpretable (with parameters) in each other if the following conditions hold:
  • 1)
    $\mathbb{A}$} and $\mathbb{B}$} are interpretable (with parameters) in each other, so ${\mathbb{A}}\simeq\Gamma({\mathbb{B}},p)$ and ${\mathbb{B}}\simeq\Delta({\mathbb{A}},q)$ for some codes $\Gamma$ and $\Delta$ and tuples of parameters $p,q$. By transitivity $\mathbb{A}$}, as well as $\mathbb{B}$}, is interpretable (with parameters) in itself, so ${\mathbb{A}}\simeq(\Gamma\circ\Delta)({\mathbb{A}},p^{\ast})$ and ${\mathbb{B}}\simeq(\Delta\circ\Gamma)({\mathbb{B}},q^{\ast})$, where $\circ$ denotes composition of interpretations and $p^{\ast}$, $q^{\ast}$ the corresponding parameters.
  • 2)
    There is a formula $\theta_{\mathbb{A}}(\bar{u},x,\bar{s})$ in the language $L({\mathbb{A}})$ such that $\theta_{\mathbb{A}}(\bar{u},x,p^{*})$ defines in $\mathbb{A}$} the isomorphism $\bar{\mu}_{\Gamma\circ\Delta}:(\Gamma\circ\Delta)({\mathbb{A}},p^{\ast})\to{\mathbb{A}}$ (more precisely, it defines the coordinate map $\mu_{\Gamma\circ\Delta}:A_{\Gamma\circ\Delta}\to A$). Similarly, there is a formula $\theta_{\mathbb{B}}(\bar{v},x,\bar{t})$ in the language $L({\mathbb{B}})$ such that $\theta_{\mathbb{B}}(\bar{v},x,q^{*})$ defines in $\mathbb{B}$} the isomorphism $\bar{\mu}_{\Delta\circ\Gamma}:(\Delta\circ\Gamma)({\mathbb{B}},q^{\ast})$ (more precisely, it defines the coordinate map $\mu_{\Delta\circ\Gamma}:B_{(\Delta\circ\Gamma)}\to B$).
Definition 2.3.
Two algebraic structures $\mathbb{A}$} and $\mathbb{B}$} are called regularly bi-interpretable in each other if the following conditions hold:
  • 1)
    $\mathbb{A}$} and $\mathbb{B}$} are regularly interpretable in each other, so ${\mathbb{A}}\simeq\Gamma({\mathbb{B}},\phi)$ and ${\mathbb{B}}\simeq\Delta({\mathbb{A}},\psi)$ for some codes $\Gamma$ and $\Delta$ and the corresponding formulas $\phi,\psi$ (without parameters). By transitivity $\mathbb{A}$}, as well as $\mathbb{B}$}, is regularly interpretable in itself, so ${\mathbb{A}}\simeq(\Gamma\circ\Delta)({\mathbb{A}},\phi^{\ast})$ and ${\mathbb{B}}\simeq(\Delta\circ\Gamma)({\mathbb{B}},\psi^{\ast})$, where $\circ$ denotes composition of interpretations and $\phi^{\ast},\psi^{\ast}$ the corresponding formulas.
  • 2)
    There is a formula $\theta(\bar{y},x,\bar{z})$ in the language of $\mathbb{A}$} such that for every tuple $p^{\ast}$ satisfying $\phi^{\ast}(\bar{z})$ in $\mathbb{A}$} the formula $\theta(\bar{y},x,p^{\ast})$ defines in $\mathbb{A}$} the isomorphism $\bar{\mu}_{\Gamma\circ\Delta}:(\Gamma\circ\Delta)({\mathbb{A}},p^{\ast})\to{\mathbb{A}}$ and there is a formula $\sigma(\bar{u},x,\bar{v})$ in the language of $\mathbb{B}$} such that for every tuple $q^{\ast}$ satisfying $\psi^{\ast}(\bar{v})$ in $\mathbb{B}$} the formula $\sigma(\bar{u},x,q^{\ast})$ defines in $\mathbb{B}$} the isomorphism $\bar{\mu}_{\Delta\circ\Gamma}:(\Delta\circ\Gamma)({\mathbb{B}},q^{\ast})\to{\mathbb{B}}$.
Algebraic structures $\mathbb{A}$} and $\mathbb{B}$} are called 0-bi-interpretable or absolulutely by-interpretable in each other if in the definition above the tuples of parameters $p$ and $q$ are empty. But 0-bi-interpretability is rather rare. Indeed, [16] if $\mathbb{A}$} and $\mathbb{B}$} are 0-bi-interpretable in each other then their groups of automorphisms are isomorphic.
2.2 Interpretability in $\mathbb{N}$ and $\mathbb{Z}$.
By Lagrange’s Four Square Theorem, every positive integer is a sum of four squares; hence $\mathbb{N}$} is a definable subset of $\mathbb{Z}$}. In particular, $\mathbb{N}$} is absolutely interpretable in $\mathbb{Z}$}. Note that every element of $\mathbb{N}$} (of $\mathbb{Z}$}) is 0-definable in $\mathbb{N}$} (in $\mathbb{Z}$}); therefore, every algebraic structure interpretable in $\mathbb{N}$} (in $\mathbb{Z}$}) with parameters is also absolutely interpretable in $\mathbb{N}$} (in $\mathbb{Z}$}). On the other hand, it is known (and easy) that $\mathbb{Z}$} is also interpretable in $\mathbb{N}$}. This can be done in a number of different ways. For example, non-negative integers $n$ can be interpreted in $\mathbb{N}$} by pairs $(0,n)$, while negative integers $m$ can be interpreted by pairs $(1,|m|)$. It is not hard to write formulas for such pairs that would interpret addition and multiplication in $\mathbb{Z}$}, so $\mathbb{N}$} and $\mathbb{Z}$} are mutually interpreted in each other. In fact, the following is true (and easy to prove; see, for example, [17, Theorem 1.5.13])
Proposition 2.4.
$\mathbb{N}$} and $\mathbb{Z}$} are absolutely bi-interpretable in each other.
From the transitivity of interpretations (see [8]), it follows that an algebraic structure $\mathcal{A}$} is interpretable (regularly interpretable) in $\mathbb{N}$} if and only if it is interpretable (regularly interpretable) in $\mathbb{Z}$}. It is convenient because many classical results are about interpretations in $\mathbb{N}$}, but in groups, interpretations with $\mathbb{Z}$} are more natural.
From the arithmetical hierarchy and its relation to Turing degrees (usually referred to as Post’s theorem), it follows that every computably enumerable subset of ${\mathbb{N}}^{n}$, $n\geq 1$, is absolutely definable in $\mathbb{N}$} (see [41, 45]). This provides a very useful tool for interpretability of algebraic structures in $\mathbb{N}$}. To make it more precise, recall that an algebraic structure ${\mathcal{A}}=\langle A;L\rangle$ in a finite language $L$ is computable or recursive (see [39, 25]) if there is a surjective map $\nu:{\mathbb{N}}\to{\mathcal{A}}$, called an enumeration of $\mathcal{A}$}, such that the full preimages with respect to $\nu$ of all the graphs of the basic operations and predicates, including the equality =, on $\mathcal{A}$} are computable in $\mathbb{N}$}. For such an enumeration $\nu$ there is a computable function $\lambda:{\mathbb{N}}\to{\mathbb{N}}$, where $\lambda(i)=j$ if $j$ is the smallest natural number such that $\nu(j)=\nu(i)$. Put ${\mathbb{N}}_{0}=\lambda({\mathbb{N}})$. Then the restriction $\nu_{0}$ of $\nu$ to ${\mathbb{N}}_{0}$ gives a bijective enumeration $\nu_{0}:{\mathbb{N}}_{0}\to A$ such that the full $\nu_{0}$-preimages of the basic operations and predicates on $A$ are computable in $\mathbb{N}$}. Since $\nu_{0}$ is a bijection, this construction gives an $L$-structure denoted by ${\mathcal{A}}_{\nu}$ on ${\mathbb{N}}_{0}$ that is isomorphic to $\mathcal{A}$} with respect to $\nu_{0}$. Since the basic operations and predicates are computable in $\mathbb{N}$}, they are definable in $\mathbb{N}$}; hence, ${\mathcal{A}}_{\nu}$ is an interpretation of $\mathcal{A}$} in $\mathbb{N}$} with the coordinate map $\nu_{0}$.
In particular, a finitely generated group $G$ with a decidable word problem is computable, since one can effectively enumerate all group words in a fixed finite set of generators of $G$, thus obtaining an enumeration $\nu:{\mathbb{N}}\to G$ that makes the group $G$ computable. Therefore, such groups are interpretable in $\mathbb{N}$}, hence in $\mathbb{Z}$}.
Similarly, every finitely generated commutative ring with a decidable equality problem is computable; hence, it is interpretable in $\mathbb{N}$} and in $\mathbb{Z}$}. It follows, for example, that a ring of Laurent polynomials ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{n},a_{n}^{-1}]$ is interpretable in $\mathbb{N}$} and in $\mathbb{Z}$}. Furthermore, it was shown in [27] that ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{n},a_{n}^{-1}]$ is bi-interpretable with $\mathbb{N}$}.
2.3 Wreath products of free abelian groups
Let $A$ and $B$ be free abelian groups with bases $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$, respectively. If not stated otherwise, we view the groups $A$ and $B$ in multiplicative notation. Denote by $G=B\,{\rm wr}\,A$ the restricted wreath product of $A$ and $B$. We fix this notation throughout the entire paper.
By construction, the group $G$ is a semidirect product $N\rtimes A$, where $N$ is the direct product $N\simeq\Pi_{a\in A}B_{a}$ of isomorphic copies $B_{a}$ of $B$, termed the base group of this wreath product, and $A$, termed the active group, acts on $N$ by shifting the copies $B_{a}$ in $N$. Namely if $f\in\Pi_{a\in A}B_{a}$ is viewed as a function $f:A\to\cup_{a\in A}B_{a}$, then the action of $a\in A$ on $f$ results in a function $f^{a}(x)=f(ax)$ for any $x\in A$. As usual, elements $g\in G$ can be written uniquely in a normal form $g=af$, where $a\in A,f\in N$, and the multiplication is given by
$a_{1}f_{1}\cdot a_{2}f_{2}=a_{1}a_{2}f_{1}^{a_{2}}f_{2}.$ (3)
Note that in this case $f^{a}=a^{-1}fa$ - the conjugation of $f$ by $a$. The groups $A$ and $B$ canonically embed into $G$ via the maps $a\in A\to a\cdot 1_{N}\in G$ and $b\in B\to f_{b}$, where $f_{b}(x)=1$ for $x\neq 1$ and $f_{b}(1)=b$. We often identify $A$ and $B$ with their images in $G$, so $B$ is identified with the copy $B_{1}$.
We list below some known properties of wreath products $G$.
  1. 1)
    The subgroup $N$ of $G$ is precisely the normal closure $N=ncl(B)$ of $B$ in $G$.
  2. 2)
    The action of $A$ on $N$ extends linearly to the action of the integer group ring ${\mathbb{Z}}A$ on $N$, so for $u=\Sigma n_{i}a_{i}\in{\mathbb{Z}}A$ and $t\in N$ $t^{u}=\Pi(t^{n_{i}})^{a_{i}}$. This makes $N$ into ${\mathbb{Z}}A$-module. Observe, that we use here the multiplicative notation for the action of ${\mathbb{Z}}A$ on $N$, which matches naturally with the conjugation notation in $G$.
  3. 3)
    The ring ${\mathbb{Z}}A$ is isomorphic to the ring of Laurent polynomials
    ${\mathbb{Z}}A\simeq{\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}],$
    so sometimes we identify ${\mathbb{Z}}A$ with $Z[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$.
  4. 4)
    ZAZA-module NN is a free ${\mathbb{Z}}A$-module with basis $b_{1},\ldots,b_{n}$.
  5. 5)
    Every element $g\in G$ can be written uniquely in the canonical (normal) form
    $g={a_{1}}^{\gamma_{1}}\ldots{a_{m}}^{\gamma_{m}}{b_{1}}^{P_{1}}\ldots{b_{n}}^{P_{n}}=\Pi_{i}a_{i}^{\gamma_{i}}\Pi_{j}b_{j}^{P_{j}},$
    where $\gamma_{i}\in{\mathbb{Z}}$ and $P_{1},\ldots,P_{n}$ are Laurent polynomials from ${\mathbb{Z}}A$. The multiplication of normal forms in $G$ is given by the following formula:
    $\Pi_{i}a_{i}^{\gamma_{i}}\Pi_{j}b_{j}^{P_{j}}\cdot\Pi_{i}a_{i}^{\beta_{i}}\Pi_{j}b_{j}^{Q_{j}}=\Pi_{i}a_{i}^{\gamma_{i}+\beta_{i}}\Pi_{j}b_{j}^{P_{j}\Pi_{i}a_{i}^{\beta_{i}}+Q_{j}}$ (4)
The tuple $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ plays a role of a ”basis” of $G$ and for $g\in G$ the tuple
$s_{p}(g)=(\gamma_{1},\ldots,\gamma_{m},P_{1},\ldots,P_{n})$
is the tuple of ”coordinates” of $g$, provided $g=\Pi_{i}a_{i}^{\gamma_{i}}\Pi_{j}b_{j}^{P_{j}}$.
For convenience, in this and similar cases, we write sometimes $p=\bar{a}\cdot\bar{b}$, where $\bar{a}=(a_{1},\ldots,a_{m})$, $\bar{b}=(b_{1},\ldots,b_{n})$ and $\cdot$ is concatenation of tuples. Respectively, we write $s_{p}(g)=\bar{\gamma}\cdot\bar{P}$ for the tuple of coordinates of $g$ relative to $p$. Using this notation we sometimes write $g={\bar{a}}^{\bar{\gamma}}\cdot{\bar{b}}^{\bar{P}}$.
Now we describe other bases of $G$ more formally.
Definition 2.5.
A tuple $q=(c_{1},\ldots,c_{m},u_{1},\ldots,u_{n})$, where $c_{i}\in G\smallsetminus N,u_{j}\in N,$ is termed a basis of $G=B\,{\rm wr}\,A$ if it satisfies the following equivalent conditions in $G$:
  1. 1)
    We fix the tuple $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$. The tuples $p$ and $q$ are automorphically equivalent in $G$, i.e., there is an automorphism of $G$ that maps $p$ to $q$ coordinate-wise.
  2. 2)
    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$, $G=C\cdot N$, $C\cap N=1$, and $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation.
  3. 3)
    Every $g\in G$ can be uniquely written in the form $g=\Pi_{i}c_{i}^{\gamma_{i}}\Pi_{j}u_{j}^{P_{j}}$, where $\gamma_{i}\in{\mathbb{Z}}$ and $p_{j}\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ and multiplication of these normal forms is given by the formula (4), where each $a_{i}$ is replaced by $c_{i}$, and each $b_{i}$ is replaced by $u_{i}$.
To verify that conditions 1), 2), and 3) above are equivalent, suppose first that an automorphism $\phi$ of $G$ maps $p$ to $q$. Then $\phi(A)=C$, so $C\simeq A$. By Lemma 2.7, item 2), $N$ is the centralizer of the commutant $G^{\prime}$ of $G$, hence $\phi(N)=N$. Clearly, $G=C\cdot N$ and $C\cap N=1$. The action of ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$ on $N$ comes from the conjugation of the elements $a_{i}$ on $N$, so $\phi$ maps this action to the action of ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ on $N$, and $N$ becomes a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation. Hence, 1) implies 2). Assume now that 2) holds; then $U=\langle u_{1},\ldots,u_{n}\rangle$ is isomorphic to $B$. Moreover, $G\simeq N\rtimes C$ and $G\simeq C\,{\rm wr}\,U$. Hence, the property 5) above holds for $G\simeq N\rtimes C$ and the formula (4) follows. This proves 3). To see that 3) implies 1) observe that if 3) holds then the coordinate-wise map $\psi_{0}:q\to p$ extends to a bijection on the normal forms
$\psi:\Pi_{i}c_{i}^{\gamma_{i}}\Pi_{j}u_{j}^{P_{j}}\to\Pi_{i}a_{i}^{\gamma_{i}}\Pi_{j}b_{j}^{P_{j}},$
which is a group homomorphism since the multiplication is given by the corresponding formulas for $q$ and for $p$ described in (4). Hence $\psi:G\to G$ is an automorphism, so 1) holds.
If $q=(c_{1},\ldots,c_{m},u_{1},\ldots,u_{m})$ is a basis of $G$ then we call the tuple $\bar{c}=(c_{1},\ldots,c_{m})$ a top basis of $G$ and $\bar{u}=(u_{1},\ldots,u_{m})$ a bottom basis.
Now we describe the members $G_{n}$ of the lower central series of $G$. Recall that $G_{1}=G$ and $G_{i+1}=[G,G_{i}]$ for $1<i\in{\mathbb{N}}$.
Denote by $\Delta$ the augmentation ideal of ${\mathbb{Z}}A$, that is the kernel of the homomorphism ${\mathbb{Z}}A\to{\mathbb{Z}}$ induced by the trivial homomorphism $A\to 1$. As an ideal $\Delta$ is generated by elements $a_{1}-1,\ldots,a_{m}-1$ of ${\mathbb{Z}}A$. To see this note that $a_{i}^{-1}-1=\frac{1-a_{i}}{a_{i}}$ and uv1=u(v1)+(u1)uv-1=u(v-1)+(u-1) for $u,v\in A$. Put $\Delta^{0}=1$ and let $\Delta^{i}$ be the $i$th-power of $\Delta$ for $i>0$, i.e., it is an ideal in ${\mathbb{Z}}A$ generated by all products of the type $(y_{1}-1)\ldots(y_{i}-1)$, where $y_{j}\in\{a_{1},\ldots,a_{m}\}$. Denote by $N^{\Delta^{i}}$ the ${\mathbb{Z}}A$-submodule of $N$ generated by all elements $u^{P}$, where $u\in N$ and $P\in\Delta^{i}$. Put $R={\mathbb{Z}}A$, $R_{i}={\mathbb{Z}}A/\Delta^{i}$, and $N_{i}=N/N^{\Delta^{i}}$. In particular, $R_{0}\simeq R$, $R_{1}\simeq{\mathbb{Z}}$.
Lemma 2.6.
The following hold in $G$:
  • 1)
    $G_{i}$ is generated as a normal subgroup in $G$ by all simple left-normed commutators of the type $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]$, where $1\leq k\leq n$, $1\leq j_{1},\ldots,j_{i-1}\leq m$;
  • 2)
    $G_{i}=N^{\Delta^{i-1}}$ for every natural $i>1$, in particular, $G_{2}=[G,G]=N^{\Delta}$;
  • 3)
    $N_{i}=N/N^{\Delta^{i}}\simeq B\otimes_{\mathbb{Z}}R_{i}$ for every natural $i\geq 1$;
  • 4)
    $G/G_{i}\simeq N_{i-1}\rtimes A\simeq(B\otimes_{\mathbb{Z}}R_{i-1})\rtimes A$ for every natural $i\geq 1$;
  • 5)
    $G_{i}/G_{i+1}$ is free abelian of finite rank for every natural $i\geq 1$;
  • 6)
    $G/G_{i+1}$ is torsion-free for every natural $i>1$.
Proof.
To prove 1) recall that $G_{i}$ is generated as a normal subgroup by all simple left-normed commutators of the form $x_{1},\ldots,x_{i}]$ where $x_{k}\in\{a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}\}$, $k=1,\ldots,i$ (this is true for any group $G$ and $x_{i}$ from any fixed generating set of $G$). Since $N$ is abelian and $G^{\prime}\leq N$ any commutator $x_{1},\ldots,x_{i}]$ where $x_{s}\in N$ for some $s>2$ is equal to 1. Therefore, $G_{i}$ is generated as a normal subgroup in $G$ by the commutators $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]$, as claimed.
To prove 2), note that
$b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]=b_{k}^{(a_{j_{1}}-1)\ldots(a_{j_{i-1}}-1)},$
which implies 2).
To see 3) observe that $N\simeq B\otimes_{Z}{\mathbb{Z}}A$ and $B\otimes_{\mathbb{Z}}R_{i}$ is a free $R_{i}$-module with basis $b_{1},\ldots,b_{n}$. The natural epimorphism ${\mathbb{Z}}A\to R_{i}={\mathbb{Z}}A/\Delta^{i}$ turns $B\otimes_{\mathbb{Z}}R_{i}$ into a ${\mathbb{Z}}A$-module. Now, the identical homomorphism $B\to B$ and the epimorphism ${\mathbb{Z}}A\to R_{i}$ give rise to an epimorphism
$\theta:B\otimes_{Z}{\mathbb{Z}}A\to B\otimes_{\mathbb{Z}}R_{i}.$
To prove 3) it suffices to note that $\ker\theta=N^{\Delta^{i}}$.
Clearly, 2) and 3) imply 4).
For 5) note that $G_{i}/G_{i+1}$ is freely generated as an abelian group by elements $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}],$ where $j_{1}\leq\ldots\leq j_{i-1},k=1,\ldots,n.$ Indeed, it follows from 1) and 2) that $G_{i}/G_{i+1}$ is isomorphic as an abelian group to $N^{\Delta^{i-1}}/N^{\Delta^{i}}$. It also follows from 2) that the images of the elements $b_{k}^{(a_{j_{1}}-1)\ldots(a_{j_{i-1}}-1)}$ in the quotient $N^{\Delta^{i-1}}/N^{\Delta^{i}}$ generate this quotient. Since the ring ${\mathbb{Z}}A$ is commutative we can always reorder factors in $(a_{j_{1}}-1)\ldots(a_{j_{i-1}}-1)$} and assume that $j_{1}\leq\ldots\leq j_{i-1}$. This set still generates the quotient as an abelian group. Moreover, there are no non-trivial relations between these elements. To see this, it suffices to notice that every polynomial $P\in{\mathbb{Z}}[a_{1},\ldots,a_{m}]$ can be uniquely decomposed as a polynomial with integer coefficients in ”new variables” $y_{i}=a_{i}-1$. Indeed, the map $a_{1}\to a_{1}-1,\ldots,a_{m}\to a_{m}-1$ extends to an endomorphism $\psi$ of ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$ (since ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$ is a free commutative ring with basis $a_{1},\ldots,a_{m}$). In fact, $\psi$ is a bijection because there is an inverse map induced by the map $a_{1}\to a_{1}+1,\ldots,a_{m}\to a_{m}+1$. It follows that $P=\psi(\psi^{-1}(P))$ is the required decomposition. It is unique since $\psi$ is an automorphism. Now it is clear that the elements $b_{k}^{(a_{j_{1}}-1)\ldots(a_{j_{i-1}}-1)},$ where $j_{1}\leq\ldots\leq j_{i-1},k=1,\ldots,n$ form a basis of $N^{\Delta^{i-1}}$ modulo $N^{\Delta^{i}}$. This proves 5).
6) follows from 5) since extensions of torsion-free groups are again torsion-free.
Lemma 2.7.
The following holds in $G$:
  • 1)
    NN is the centralizer $C_{G}(g)$ of any non-trivial $g\in N$;
  • 2)
    NN is the centralizer $C_{G}(G^{\prime})$ of the commutant $G^{\prime}=[G,G]$;
  • 3)
    AA is the centrailizer $C_{G}(a)$ of any non-trivial $a\in A$.
Proof.
For any non-trivial $g\in N$, every element in $N$ commutes with $g$ and every element in $A$ does not commute with $g$. Since elery element in $G$ is a product of an element in $N$ and an element in $A$, this gives 1) and 2). Since $A$ is abelian, for any $a\in A$, $A\subseteq C_{G}(a)$. And $a$ does not commute with any non-trivial $g\in N$, this gives 3).
3 Interpretability of $G$ and $\mathbb{Z}$} in each other
3.1 Interpretability of $G$ in $\mathbb{Z}$}
The group $G$ is a finitely generated metabelian group; hence, it has a decidable word problem (see, for example, [5]). Therefore, as we mentioned in Section 2.2, $G$ is absolutely interpretable in $\mathbb{Z}$}. However, we describe a particular interpretation here, as we will use it in the sequel.
We show the interpretability of $G$ in $\mathbb{Z}$ in two steps. First, we prove that the ring
$R=\mathbb{Z}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$
is interpretable in $\mathbb{Z}$. Then, using a fixed basis of $G$, we interpret $G$ in $\mathbb{Z}$ via the coordinate representations of its elements with respect to this basis.
To see that $R$ is absolutely interpretable in $\mathbb{Z}$}, note that $R$ is computable (since the word problem in $R$ is decidable); hence, there exists a computable injective function $\nu:R\to{\mathbb{N}}$ such that the set $\nu(R)$ is computable in $\mathbb{N}$} and the images under the map $\nu$ of the ring operations of the ring $R$ are computable in $\mathbb{N}$}, i.e., the following operations on $\nu(R)$ (here $Q_{i}\in R$, $i=1,2,3$) are computable in $\mathbb{N}$}:
$k_{1}\oplus k_{2}=k_{3}\Longleftrightarrow\wedge_{i=1}^{3}(k_{i}=\nu(Q_{i}))\wedge(Q_{1}+Q_{2}=Q_{3}),$
$k_{1}\odot k_{2}=k_{3}\Longleftrightarrow\wedge_{i=1}^{3}(k_{i}=\nu(Q_{i}))\wedge(Q_{1}\cdot Q_{2}=Q_{3}).$
Again, this is a general fact about computable algebraic structures. Nevertheless, it is convenient to give a sketch of a particular such enumeration $\nu:R\to{\mathbb{N}}$. Every polynomial $P\in{\mathbb{Z}}[a_{1},\ldots,a_{n}]$ can be uniquely presented as an integer linear combination of pair-wise distinct monomials on commuting variables $a_{1},\ldots,a_{n}$:
$P=\Sigma_{i=1}^{d}\gamma_{i}a_{1}^{\alpha_{i1}}\ldots a_{n}^{\alpha_{in}},$
where $0\neq\gamma_{i}\in{\mathbb{Z}}$ and $d,\alpha_{i}\in{\mathbb{N}}$. Hence, the polynomial $P$ is uniquely presented by a tuple
$u=(\gamma_{1},\alpha_{11},\ldots\alpha_{1n},\ldots,\gamma_{d},\alpha_{d1},\ldots,\alpha_{dn}).$ (5)
If $Q\in R$, then $Q$ can be uniquely presented in the form $Q=\frac{P}{{\bar{a}}^{\bar{\beta}}}$ for some $P\in{\mathbb{Z}}[a_{1},\ldots,a_{n}]$ and some monomial ${\bar{a}}^{\bar{\beta}}=a_{1}^{\beta_{1}}\ldots a_{n}^{\beta_{n}},\beta_{i}\in{\mathbb{N}}$, such that $gcd(P,{\bar{a}}^{\bar{\beta}})=1$. It follows that $Q$ can be uniquely presented by a pair of tuples $(u_{Q},v_{Q})$, where $u_{Q}=u,v_{Q}=\bar{\beta}$. Fix an arbitrary computable bijection
$\tau:\bigcup_{i\in{\mathbb{N}}}{\mathbb{Z}}^{i}\to{\mathbb{N}}$
which enumerates all finite tuples of integers. Then $Q$ is uniquely presented by the pair $(\tau(u_{Q}),\tau(v_{Q}))\in{\mathbb{N}}^{2}$ and the set of all such pairs is a computable subset of ${\mathbb{N}}^{2}$. For a fixed computable bijection $\tau_{2}:{\mathbb{N}}^{2}\to{\mathbb{N}}$, put
$\nu(Q)=\tau_{2}((\tau(u_{Q}),\tau(v_{Q}))).$
By construction, the subset $\nu(R)$ is computable in $\mathbb{N}$}, and given a number $k\in\nu(R)$, one can algorithmically find the corresponding Laurent polynomial $Q$ such that $k=\nu(Q)$. Then it is easy to see that the operations $\oplus$ and $\odot$ on $\nu(R)$ are computable in $\mathbb{N}$}. To finish the proof, recall from Section 2.2 that all the computable operations or predicates on $\mathbb{N}$} are definable in $\mathbb{N}$}, and $\mathbb{N}$} is definable in $\mathbb{Z}$}. This shows that $R$ is absolutely interpretable in $\mathbb{Z}$}.
Now, for the second step, we use notation from Section 2.3. Let $p=\bar{a}\cdot\bar{b}$ be a fixed basis of $G$. Then every element $g\in G$ can be uniquely represented as
$g={a_{1}}^{\gamma_{1}}\ldots{a_{m}}^{\gamma_{m}}{b_{1}}^{P_{1}}\ldots{b_{n}}^{P_{n}},$
so it can be uniquely represented by the tuple of its coordinates with respect to $p$
$s_{p}(g)=(\gamma_{1},\ldots\gamma_{m},P_{1},\ldots,P_{n}).$ (6)
Hence, by a tuple of integers
$s_{p,\nu}(g)=(\gamma_{1},\ldots\gamma_{m},\nu(P_{1}),\ldots,\nu(P_{n})).$ (7)
As mentioned above, the set $\nu(R)$ is definable in $\mathbb{Z}$}; therefore, the set
$S_{p,\nu}=\{s_{p,\nu}(g)\mid g\in G\}\subset{\mathbb{Z}}^{m+n}$
is also definable in $\mathbb{Z}$}. Multiplication of normal forms in $G$ with respect to the basis $p$ is given by the formula (4) from Section 2.3, which shows that the tuple of coordinates $s_{p}(gh)$ of the product of two elements $g,h\in G$ can be computed from the coordinates $s_{p}(g)$ and $s_{p}(h)$. The same holds for tuples $s_{p,\nu}(g)$, $s_{p,\nu}(h)$, and $s_{p,\nu}(gh)$. Hence, there is a formula $\Phi(\bar{x}_{1},\bar{x}_{2},\bar{x}_{3})$ of arithmetic such that for any $\bar{s}_{1},\bar{s}_{2},\bar{s}_{3}\in S_{p,\nu}$, the following holds:
${\mathbb{Z}}\models\Phi(\bar{s}_{1},\bar{s}_{2},\bar{s}_{3})\Longleftrightarrow s_{1}=s_{p,\nu}(g)\wedge s_{2}=s_{p,\nu}(h)\wedge s_{3}=s_{p,\nu}(gh)$
for some $g,h\in G$. Note that in any group, the inversion operation and the identity element are absolutely definable from the multiplication. Hence, the construction above gives an interpretation of $G$ in $\mathbb{Z}$, which we denote by $\Gamma$, so $G\simeq\Gamma(\mathbb{Z})$, with the coordinate map $\mu:s_{p,\nu}(g)\mapsto g$.
3.2 Interpretability of $\mathbb{Z}$} in $G$
In this section, we show how to interpret $\mathbb{Z}$} in $G$. Throughout, we adopt the notation introduced in Section 2.3.
Lemma 3.1.
  • 1)
    The subgroup $N$ is 0-definable in $G$ by the formula
    $\phi_{N}(x)=\forall y\forall z[x,[y,z]]=1.$
  • 2)
    The subgroup $A$ is definable in $G$ with parameters by the formula
    $\phi_{A}(x)=[x,a]=1,$
    where $1\neq a$ is an arbitrary non-trivial element of $A$.
Proof.
It follows immediately from Lemma 2.7. ∎
Lemma 3.2.
Let $1\neq a\in A$ and $g\in G$. Then the following conditions are equivalent:
  1. 1)
    $g$ belongs to the cyclic subgroup $\langle a\rangle$,
  2. 2)
    $[g,a]=1$ and there is $z\in N$ such that $b_{1},g]=[z,a]$,
  3. 3)
    $[g,a]=1$ and for any $u\in N$ there exists $z\in N$ such that $[u,g]=[z,a]$.
Proof.
1) $\Longrightarrow$ 3). Let $g\in\langle a\rangle$, so $g=a^{\gamma}$ for some $\gamma\in{\mathbb{Z}}$. Clearly $[g,a]=1$. Fix $u\in N$. If $\gamma>0$, then take
$z=u^{a^{\gamma-1}+a^{\gamma-2}+\ldots+a+1}.$
In this case
$u,g]=u^{-1}u^{a^{\gamma}}=u^{a^{\gamma}-1}=u^{({a^{\gamma-1}+a^{\gamma-2}+\ldots+1})(a-1)}=[z,a],$
as claimed. If $\gamma<0$ then take
$z=u^{\frac{-(a^{|\gamma|-1}+a^{|\gamma|-2}+\ldots+1)}{a^{|\gamma|}}}.$
The case $\gamma=0$ is obvious.
Clearly, 3) $\Longrightarrow$ 2). To prove the lemma, it suffices to show that 2) $\Longrightarrow$ 1). Let $g\in A$ such that
$b_{1},g]=[z,a]$ (8)
for some $g\in A$, $g\neq 1$ and $z\in N$. We need to show that $g=a^{\gamma}$ for some $\gamma\in{\mathbb{Z}}$.
Note first that in a free abelian group $A$ there is an automorphism $\theta\in\mathrm{Aut}A$ that maps $a$ to $a_{1}^{\beta}$ for some positive $\beta\in{\mathbb{N}}$. Replacing the basis $\{a_{1},\ldots,a_{m}\}$ of $A$ with basis $\{\theta(a_{1}),\dots,\theta(a_{m})\}$, if necessary, we may assume that $a=a_{1}^{\beta}$.
If $z=1$ then $g=1=a^{0}$. Suppose $z\neq 1$ and write $z$ as $z=\prod_{i}b_{i}^{Q_{i}}$ for some Laurent polynomials $Q_{i}\in{\mathbb{Z}}A$. Then the equality (8) above takes the form
$b_{1}^{g-1}=\Pi_{i}b_{i}^{Q_{i}(a_{1}^{\beta}-1)}.$
Since $b_{1},\ldots,b_{n}$ is a basis of the free module $N$, one gets $Q_{i}=0$ for all $i>1$. Hence
$g-1=Q_{1}(a_{1}^{\beta}-1)$
in the ring ${\mathbb{Z}}A$. Since $g\in A$ it is a product $g=\Pi_{i=1}^{m}a_{i}^{\gamma_{i}}$ for some $\gamma_{i}\in{\mathbb{Z}}$, we can write $g=\frac{g_{1}}{g_{2}}$, where $g_{1}$ is a product of all $a_{i}^{\gamma_{i}}$ with $\gamma_{i}>0$ and $g_{2}$ is a product of all $a_{i}^{|\gamma_{i}|}$ with $\gamma_{i}<0$. Note that $g_{1}$ and $g_{2}$ do not have common variables. Now,
$\frac{g_{1}-g_{2}}{g_{2}}=Q_{1}(a_{1}^{\beta}-1)$
and therefore $Q_{1}=\frac{P_{1}}{g_{2}}$, where $P_{1}$ is an ordinary polynomial from ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$. Hence, we have an equality
$g_{1}-g_{2}=P_{1}(a_{1}^{\beta}-1)$ (9)
in the ring of polynomials ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$. Consider a monomial order in monomials of ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$, for example, the deglex order where monomials are compared first by their degree and if they have the same degree then with respect to a lexicographical ordering. Write $P_{1}$ as a sum of terms $M_{j}$ (i.e., monomials with integer coefficients)
$P_{1}=M_{1}+\ldots+M_{k}$
where $M_{1}>M_{2}>\ldots M_{k}$ in the deglex order. Then the equation (9) becomes
$g_{1}-g_{2}=M_{1}a_{1}^{\beta}-M_{1}+M_{2}a_{1}^{\beta}-M_{2}+\ldots+M_{k}a_{1}^{\beta}-M_{k}.$ (10)
Clearly, $M_{1}a_{1}^{\beta}$ is the leading term on the right and $M_{k}$ is the smallest one, so they both cannot cancel out on the right-hand side. Then, assuming $g_{1}>g_{2}$ we have $g_{1}=M_{1}a_{1}^{\beta}$ and $g_{2}=M_{k}$. Observe, that $g_{1}$ contains $a_{1}$ and $g_{2}$ does not. Now, the equality (10), after canceling equal terms, turns into
$M_{1}-g_{2}=M_{2}a_{1}^{\beta}-M_{2}+\ldots-M_{k}.$
which again has the form (10). By induction, $M_{i}=M_{i+1}a_{1}^{\beta}$ for all $i=1,\ldots,k-1$. In particular, $g_{2}$ divides $g_{1}$. Since $g_{1}$ and $g_{2}$ do not have common divisors besides $1$ and 1-1, we deduce that $g_{2}=\pm 1$, and from $g=\frac{g_{1}}{g_{2}}$ we have $g_{2}=1$. It follows that $P_{1}=a_{1}^{\beta(k-1)}+\ldots a_{1}^{\beta}+1$ and $g=g_{1}=a^{k\beta}$ for some $k>0$. However, if we assume above that $g_{2}$ contains $a_{1}$. Then a similar argument shows that in this case $g_{1}=1$ and $g_{2}=a_{1}^{k\beta}$ for some $k>0$, so $g=a_{1}^{-k\beta}$, as claimed.
The two distinct yet equivalent conditions (2) and (3) in Lemma 3.2 make it possible to define the cyclic subgroup $\langle a\rangle$ of $G$ for any $a\in A,a\neq 1,$ using different formulas. This flexibility is particularly useful when investigating the decidability of equations in $G$ or addressing other problems where the complexity of the defining formulas plays a crucial role. We summarize this observation in the following corollary.
Corollary 3.3.
The following holds in $G$:
  1. 1.
    There is a $\forall\exists$ positive formula $Cyc_{A}(x,y)$ of the group theory language such that for any $a\in A$ the formula $Cyc_{A}(a,y)$ defines the cyclic subgroup $\langle a\rangle$ in $G$.
  2. 2.
    There is a Diophantine formula $Cyc^{\prime}_{A}(x,y,b_{1})$ with the parameter $b_{1}$ such that for any $1\neq a\in A$ the formula $Cyc_{A}(a,y,b_{1})$ defines the cyclic subgroup $\langle a\rangle$ in $G$.
Proof.
It follows directly from Lemma 3.2 and Lemma 3.1, the latter is needed to define NN in $G$ without parameters. ∎
Lemma 3.4.
There is a fomula $Base_{top}(x_{1},\ldots,x_{m})$ of group language (without parameters) such that for any $c_{1},\ldots,c_{m}\in G$ $Base_{top}(c_{1},\ldots,c_{m})$ holds in $G$ if and only if the tuples $(c_{1},\ldots,c_{m})$ and $(a_{1},\ldots,a_{m})$ are automorphically equivalent in $G$, i.e., $(c_{1},\ldots,c_{m})$ is a top basis of $G$.
Proof.
By Corollary 3.3 the formula $Cyc_{A}(a_{i},y)$ defines in $G$ the cyclic subgroup $\langle a_{i}\rangle$. It follows that the tuple $(a_{1},\ldots,a_{m})$ satisfies all the conditions below:
  1. 1)
    $a_{i}\in G\smallsetminus N$ and $a_{i},a_{j}]=1$ for all $1\leq i,j\leq m$.
  2. 2)
    The set $\{g\in G\mid G\models Cyc_{A}(a_{i},g)\}$ is an abelian subgroup of $G$, containing $a_{i}\neq 1$. Denote this subgroup by $H_{i}$.
  3. 3)
    $H_{i},H_{j}]=1$, so $H=H_{1}\cdot\ldots\cdot H_{m}$ is an abelian group.
  4. 4)
    $H_{i}\cap H_{j_{1}}\cdot\ldots\cdot H_{j_{m-1}}=1$ for any $1\leq j_{k}\leq m,j_{k}\neq i$.
  5. 5)
    $H\cap N=1$.
  6. 6)
    $G=HN$.
It is clear that each one of these conditions can be written by a formula of group language on elements $a_{1},\ldots,a_{m}$. Replacing in this formula every occurence of the element $a_{i}$ by a variable $x_{i}$ we obtain a formula of group language that we denote by $Base_{top}(x_{1},\ldots,x_{m})$.
Observe, that $G\models Base_{top}(a_{1},\ldots,a_{m})$, in fact, this it true for every basis $(a_{1}^{\prime},\ldots,a_{m}^{\prime})$ of $A$, because from the beginning $(a_{1},\ldots,a_{m})$ was chosen as an arbitrary basis of $A$. Suppose that $G\models Base_{top}(c_{1},\ldots,c_{m})$ for some elements $c_{1},\ldots,c_{m}\in G$. Then for every $c_{i}$ the formula $Cyc_{A}(c_{i},y)$ defines in $G$ an abelian subgroup $C_{i}$, which contains a non-trivial element $c_{i}$. These subgroups satisfy all the conditions 1)-6), hence
  1. 7)
    $C=C_{1}\ldots C_{m}$ is an abelian subgroup of $G$.
  2. 8)
    $C\simeq C_{1}\times\ldots C_{m}\simeq A$.
  3. 9)
    $G=N\rtimes C$.
From 7) it follows that $C$ is a free abelian of rank $m$, so every group $C_{i}$ is cyclic. Denote by $d_{i}u_{i}$, where $d_{i}\in A,u_{i}\in N$, a generator of $C_{i}$. It is easy to see that $d_{1},\dots,d_{m}$ form a basis of $A$. Therefore, there is an automorphism $\psi\in Aut(A)$ such that $\psi(a_{i})=d_{i},i=1,\ldots,m$. We claim that the automorphism $\psi$ extends to an automorphism $\hat{\psi}$ of $G$. Clearly, $\psi$ extends linearly to an automorphism $\lambda_{\psi}$ of ${\mathbb{Z}}A$. Now, $\lambda_{\psi}$ gives rise to a semilinear bijection $\tilde{\psi}$ of the ${\mathbb{Z}}A$-module $N$, where for $u=\Pi b_{i}^{P_{i}}\in N,P_{i}\in{\mathbb{Z}}A$,
${\tilde{\psi}}(u)=\Pi b_{i}^{\lambda_{\psi}(P_{i})}.$
Now one can define an a map $\hat{\psi}$ on $G$ by $\hat{\psi}(au)=\psi(a){\tilde{\psi}}(u)$, where $a\in A,u\in N$. Clearly, $\tilde{\psi}$} is a bijection on $G$. Checking that it is a homomorphism of $G$ is straightforward if using the formula (3) of the multiplication in $G$.
The isomorphism ${\tilde{\psi}}^{-1}$ maps $d_{i}u_{i}$ to $a_{i}v_{i}$, where $v_{i}={\tilde{\psi}}^{-1}(u_{i})\in N$, so $(a_{1}v_{1},\ldots,a_{m}v_{m})$ forms a basis of a free abelian group $A^{\prime}$, which is isopmorphic to $A$ via a map $\theta:a_{i}\to a_{i}v_{i},i=1,\ldots,n$. Note, that the elements $a_{i}v_{i}$ act by conjugation on $N$ precisely the same way as elements $a_{i}$, hence the map $au\to\theta(a)u$, for $a\in A,u\in N$ is an isomorphism of $G$. This shows that the tuples $(a_{1},\ldots,a_{m})$, $(a_{1},v_{1},\ldots,a_{m}v_{m})$, and $(d_{1}u_{1},\ldots,d_{m}u_{m})$ are automorphically equivalent in $G$.
Recall that $c_{i}\in C_{i}=\langle d_{i}u_{i}\rangle$, so $c_{i}=(d_{i}u_{i})^{\gamma_{i}}$, for some $\gamma_{i}\in{\mathbb{Z}}$. Now we apply Lemma 3.2 where the basis $a_{1},\ldots,a_{m}$, hence the group $A$, is replaced with the basis $(d_{1}u_{1},\ldots,d_{m}u_{m})$ and the group $A^{\prime}$, and the element $a$ is replaced by $c_{i}\in A^{\prime}$. This gives us that the formula $Cyc(c_{i},y)$ defines in $G$ the cyclic subgroup $\langle c_{i}\rangle$ which is equal to $C_{i}$. Therefore $c_{i}=(d_{i}u_{i})^{\pm 1}$, so there is an automorphism of $G$ that maps $c_{i}$ to $d_{i}u_{i}$ for $i=1,\ldots,m$. This implies that the tuples $(c_{1},\ldots,c_{m})$ and $(a_{1},\ldots,a_{m})$ are automorphically equivalent in $G$, as required.
Lemma 3.5.
Let $1\neq a\in A$ and $k,\ell\in{\mathbb{Z}}$. Then the following conditions are equivalent:
  1. 1)
    $k$ divides $\ell$,
  2. 2)
    there is $z\in N$ such that $b_{1},a^{k}]=[z,a^{\ell}]$,
  3. 3)
    for any $u\in N$ there exists $z\in N$ such that $u,a^{k}]=[z,a^{\ell}].$
Proof.
The argument is similar to the one in Lemma 3.2. ∎
Corollary 3.6.
The following holds in $G$:
  1. 1.
    There exists an $\forall\exists$ and positive formula $Div_{A}(x,y_{1},y_{2})$ of the group theory language such that for any $a\in A$ and any $k,\ell\in{\mathbb{Z}}$, the formula $Div_{A}(a,a^{k},a^{\ell})$ holds in $G$ if and only if $k$ divides $\ell$.
  2. 2.
    For any $1\neq a\in A$ and any $k,\ell\in{\mathbb{Z}}$ a system of equations
    $z,b_{1}]=1\wedge[b_{1},a^{k}]=[z,a^{\ell}]$ (11)
    with parameters $a,b_{1}$ has a solution $z$ in $G$ if and only if $k$ divides $\ell$.
Proof.
The statement 2) follows from item 3) in Lemma 3.5, and 3) follows from item 2) in Lemma 3.5. ∎
Proposition 3.7.
Let $a\in A,a\neq 1$. Then the ring $\mathbb{Z}$} is interpretable in $G$ with the parameter $a$.
Proof.
We interpret $\mathbb{Z}$} in $G$ on the cyclic subgroup $\langle a\rangle=\{a^{k}\mid k\in{\mathbb{Z}}\}$. By Corollary 3.3 the set $Z_{a}=\langle a\rangle$ is definable in $G$ with the parameter $a$. We interpret addition $\oplus$ on $Z_{a}$ by the multiplication in $G$, since $a^{k}\cdot a^{\ell}=a^{k+\ell}$ the map $\mu_{a}:a^{k}\to k$ gives rise to an isomorphism $\mu_{a}:\langle Z_{a};\oplus\rangle\to\langle{\mathbb{Z}};+\rangle$.
Next, we interpret the integer division $\mid$ in $Z_{a}$ by a binary predicate $\mid^{*}$ which is defined either by the formula $Div_{A}(a,y_{1},y_{2})$ of Corollary 3.6 or by the system of equations (11) of Corollary 3.6. The map $\mu_{a}$ gives an isomorphism of the structures
$\mu_{a}:\langle Z_{a};\oplus,\mid^{*}\rangle\to\langle{\mathbb{Z}};+,\mid\rangle,$
so we interpreted $\langle{\mathbb{Z}};+,\mid\rangle$ in $G$.
Finally, there is an argument due to Robinson (see [40]) that shows that the integer multiplication $\cdot$ is definable in $\langle{\mathbb{Z}};+,\mid\rangle$ by some formula $Mult(x_{1},x_{2},x_{3})$. Using formulas for $\oplus$ and $\mid^{*}$ one can rewrite the formula $Mult(x_{1},x_{2},x_{3})$ into a formula $M_{a}(x_{1},x_{2},x_{3},a)$ with parameter $a$ which defines a multiplication $\otimes$ on $Z_{a}$ in such a way that the map $\mu_{a}$ gives an isomorphism
$\mu_{a}:\langle Z_{a};\oplus,\otimes\rangle\to\langle{\mathbb{Z}};+,\cdot\rangle.$
This gives an interpretation ${\mathbb{Z}}_{a}=\langle Z_{a};\oplus,\otimes\rangle$ of $\mathbb{Z}$} in $G$ with the parameter $a$. ∎
Remark 3.8.
We want to mention the following about the interpretations of $\mathbb{Z}$} in $G$:
  1. 1.
    For any $a\in A,a\neq 1$, the interpretation ${\mathbb{Z}}_{a}$ is injective, that is, we use the standard equality == from $G$ to define the equality in ${\mathbb{Z}}_{a}$. This might be useful for applications, in particular, in studying the elimination of imaginaries.
  2. 2.
    the structure $\langle{\mathbb{Z}};+,\mid\rangle$ is interpretable in $G$ by Diophantine formulas.
Now we show that the isomorphisms ${\mathbb{Z}}_{a}\to{\mathbb{Z}}_{d}$ are uniformly definable in $G$ with parameters $a,d\in A$.
Lemma 3.9.
Let $a,d\in A,a,d\neq 1$ and $k,l\in{\mathbb{Z}}$. Then $k=l$ if and only if for some nontrivial $u\in N$ (equivalently, every $u\in N$) the following holds:
$u,a,d^{k}]=[u,a^{\ell},d]\ \ mod\ G_{4},$
where $G_{4}$ is the fourth term of the lower central series of $G$.
Proof.
The ternary left-normed commutator $x_{1},x_{2},x_{3}]$ defines a map $N\times G\times G\to G_{3}/G_{4}$ which is linear in every variable. If $u\in N,u\neq 1$, then $u,a,d]\neq 1$. Now $u,a,d^{k}]=[u,a,d]^{k}\ \ mod\ G_{4}$ and $u,a^{\ell},d]=[u,a,d]^{\ell}\ \ mod\ G_{4}$. Hence
$u,a,d^{k}]=[u,a^{\ell},d]\ \ mod\ G_{4},$
implies that
$u,a,d]^{k}=[u,a,d]^{\ell}\ \ mod\ G_{4},$
and therefore $k=\ell$ since $G/G_{4}$ is torsion-free by Lemma 2.6.
Corollary 3.10.
There is a formula $Iso_{\mathbb{Z}}(x_{1},x_{2},y_{1},y_{2})$ in the group theory language such that for any $a,d\in A,a,d\neq 1,$ the formula $Iso_{\mathbb{Z}}(x_{1},x_{2},a,d)$ with parameters $a,d$ defines in $G$ the graph of the isomorphism ${\mathbb{Z}}_{a}\to{\mathbb{Z}}_{d}$.
Remark 3.11.
Corollary 3.10 allows one to fix an arbitrary $a\in A,a\neq 1$, make some computations with integers in ${\mathbb{Z}}_{a}$ and then transfer the result through $Iso_{\mathbb{Z}}(x_{1},x_{2},a,d)$ into ${\mathbb{Z}}_{d}$. For example, the set
$\{(a^{k},d^{2^{k}})\mid k\in{\mathbb{N}}\}$
is uniformly definable in $G$ with parameters $a,d\in A,a,d\neq 1$.
Corollary 3.12.
There is a formula $Prod_{A,m}(x_{1},x_{2},\bar{y},\bar{z})$ of the group language, here $\bar{y},\bar{z}$ are tuples of variables of length $m$, such that for any $a\neq 1,d_{1},\ldots,d_{m}\in A$, $a^{\gamma_{1}},\ldots,a^{\gamma_{m}}\in{\mathbb{Z}}_{a}$, and $g\in G$ the following holds:
$G\models Prod_{A}(g,a,\bar{d},{\bar{a}}^{\bar{\gamma}})\Longleftrightarrow g=d_{1}^{\gamma_{1}}\ldots d_{m}^{\gamma_{m}}$
where $\bar{d}=(d_{1},\ldots,d_{m}),{\bar{a}}^{\bar{\gamma}}=(a^{\gamma_{1}},\ldots,a^{\gamma_{m}})$.
In particular, the formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})=Prod_{A,2}(x_{1},x_{2},\bar{y},\bar{z})$ defines in $G$ the $\mathbb{Z}$}-exponentiation on $A$, that is for any $a,d\neq 1\in A$ and any $\gamma\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
$G\models Prod_{A}(g,a,d,a^{\gamma})\Longleftrightarrow g=d^{\gamma}.$
Corollary 3.13.
There is a formula $Base_{A}(x_{1},x_{2},\bar{y},\bar{z})$ of the group language such that for any $d_{1},\ldots,d_{m}\in A$
$G\models Base_{A}(d_{1},\ldots,d_{m})\Longleftrightarrow(d_{1},\ldots,d_{m})\text{ is a basis of}A$
4 Definability of $\mathbb{Z}$}- and ${\mathbb{Z}}A$-exponentiations in $G$
4.1 Definability of $\mathbb{Z}$}-exponentiations in $G$
We first prove some auxiliary results.
Lemma 4.1.
If $a\in A,a\neq 1,u\in N,u\neq 1,$ and $k\in\mathbb{Z}$. Then the following holds:
  1. 1)
    $(au)^{k}=a^{k}w$, where $w=u^{a^{k-1}+a^{k-2}+\ldots+1}$
  2. 2)
    this $w=u^{a^{k-1}+a^{k-2}+\ldots+1}$ is the unique solution of the equation $u,a^{k}]=[w,a]$.
Proof.
Let $k>0$. Then by induction
$(au)^{k}=(au)^{k-1}au=a^{k-1}u^{a^{k-2}+\ldots+1}au=a^{k}u^{a^{k-1}+\ldots+1},$
which proves 1). To see 2) rewrite the equality $u,a^{k}]=[w,a]$ in the form $u^{a^{k}-1}=w^{a-1}$. Dividing $a^{k}-1$ by a1a-1 one gets a unique solution $w=u^{a^{k-1}+a^{k-2}+\ldots+1}$.
Corollary 4.2.
There is a formula $Exp_{mix}(x_{1},x_{2},x_{3},y)$ of group theory language such that for any $a\neq 1\in A$, any $u\neq 1\in N$, and any $k\in{\mathbb{Z}}$ for any $g\in G$ the following holds:
$G\models Exp_{mix}(g,u,a^{k},a)\Longleftrightarrow g=(au)^{k}.$
Proof.
It follows from Lemma 4.1 and the results of the previous section.
Lemma 4.3.
Let $u\in N$, $k\in\mathbb{Z}$. Then for every $a\in A$ there exists $v\in N$ such that
$(au)^{k}=a^{k}u^{k}[v,a].$ (12)
Moreover, $w=u^{k}$ is the only element in $N$ such that for every $a\in A$ the equation
$(au)^{k}=a^{k}w[v,a]$ (13)
has a solution $v\in N$.
Proof.
We prove first, by induction on $k$, that
$u^{k}a=au^{k}[v,a]$ (14)
for some $v\in N$. For $k=1$ one has $ua=au[u,a]$. By induction
$u^{k+1}a=u^{k}ua=u^{k}au[u,a]=au^{k}[v_{1},a]u[u,a]=au^{k+1}[v_{1},a][u,a]=au^{k+1}[v_{1}u,a],$
and (14) follows.
Now we prove (12) by induction on $k$ as follows (here an element $v_{2}\in N$ appears as the result of the induction step and an element $v_{3}$ as an application of the (14)):
$(au)^{k+1}=(au)^{k}au=a^{k}u^{k}[v_{2},a]au=a^{k}u^{k}a[v_{2},a][[v_{2},a],a]u=$
$a^{k}au^{k}[v_{3},a][v_{2},a][[v_{2},a],a]u=a^{k+1}u^{k+1}[v_{3}v_{2}[v_{2},a],a]=a^{k+1}u^{k+1}[v,a],$
where $v=v_{3}v_{2}[v_{2},a],a]$.
For the uniqueness, let $u\in N,w\in G$. If for any $a\in A$, there exists $v_{1}\in N$ such that
$(au)^{k}=a^{k}w[v_{1},a].$
Then by (12) there exists $v\in N$ such that
$a^{k}u^{k}[v,a]=a^{k}w[v_{1},a].$
Hence
$w^{-1}u^{k}=[v_{1},a][v^{-1},a]=[v_{1}v^{-1},a].$
Since $N$ is a free ${\mathbb{Z}}A$-module, one gets $w^{-1}u^{k}=1.$ Indeed, let $w^{-1}u^{k}=\Pi_{i}b_{i}^{P_{i}}$ and $v_{1}v^{-1}=\Pi_{i}b_{i}^{Q_{i}}$ for some $P_{i},Q_{i}\in{\mathbb{Z}}A$. Then for every $i$ $P_{i}=Q_{i}(a-1)$. It follows that a1a-1 divides $Q_{i}$ for every $a\in A$. In particular, every polynomial $a_{1}^{\ell}-1,\ell>0$ divides every $P_{i}$. Hence, $P_{i}=0$ for all $i$, so $w^{-1}u^{k}=1$, as claimed. ∎
Corollary 4.4.
There is a formula $Exp_{N}(x_{1},x_{2},x_{3},y)$ of group theory language such that for any $a\neq 1\in A$, any $u\neq 1\in N$, and any $k\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
$G\models Exp_{N}(g,u,a^{k},a)\Longleftrightarrow g=u^{k}.$
Theorem 4.5.
There is a formula $Exp_{G}(x_{1},x_{2},x_{3},y)$ of the group theory language such that for any $a\neq 1\in A,g,h\in G,k\in\mathbb{Z}$
$G\models Exp_{G}(g,h,a^{k},a)\Longleftrightarrow g^{k}=h.$
Proof.
It follows from Corollaries 4, 6, and 7. ∎
4.2 Interpretation of the ${\mathbb{Z}}A$-action on $N$ in $G$
In this section, we show that the standard action of ${\mathbb{Z}}A$ on $N$ is interpretable in $G$ with parameters $a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}$. We fix an element $a=a_{1}\in A$ and use the interpretation ${\mathbb{Z}}_{a}$ of the ring of integers $\mathbb{Z}$} in $G$.
For a tuple $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ denote by $\lambda_{\bar{\alpha}}$ the homomorphim $\lambda_{\bar{\alpha}}:{\mathbb{Z}}[a_{1},\ldots,a_{m}]\to\mathbb{Z}$ such that $a_{i}\to\alpha_{i},i=1,\ldots,m$. The kernel $I_{\bar{\alpha}}$ of $\lambda_{\bar{\alpha}}$ is the ideal generated in ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$ by $\{a_{1}-\alpha_{1},\ldots,a_{m}-\alpha_{m}\}$. Notice, that for every polynomial $P=P(a_{1},\ldots,a_{m})\in{\mathbb{Z}}[a_{1},\ldots,a_{m}]$ one has $\lambda_{\bar{\alpha}}(P)=P(\alpha_{1},\ldots,\alpha_{m})$, so
$P(a_{1},\ldots,a_{m})=P(\alpha_{1},\ldots,\alpha_{m})+\Sigma_{i=1}^{m}(a_{i}-\alpha_{i})f_{i},$
for some $f_{i}\in{\mathbb{Z}}[a_{1},\ldots,a_{m}]$.
Let $C$ and $D$ be rings and $\Lambda$ a set of homomorphisms from $C$ into $D$. Recall that $C$ is discriminated into $D$ by a set $\Lambda$ if for any finite subset $C_{0}\subseteq C$ there is a homomorphism $\lambda\in\Lambda$ which is injective on $C_{0}$.
Lemma 4.6.
The ring ${\mathbb{Z}}[a_{1},\ldots,a_{m}]$ is discriminated into $\mathbb{Z}$} by the set of homomorphisms $\{\lambda_{\bar{\alpha}}\mid\bar{\alpha}\in{\mathbb{Z}}^{m}\}$.
Proof.
By induction on $m$. ∎
Denote by $(N)^{I_{\bar{\alpha}}}$ the submodule of the ${\mathbb{Z}}A$-module $N$ generated by $N^{I_{\bar{\alpha}}}$. Observe, that every element $u\in N^{I_{\bar{\alpha}}}$, as well as in $(N)^{I_{\bar{\alpha}}}$ admits a decomposition of the type $u=\Pi_{i}b_{i}^{U_{i}}$, where $U_{i}\in I_{\bar{\alpha}}$. As an abelian group, $(N)^{I_{\bar{\alpha}}}$ is generated by the set $\{g^{Q}\mid g\in N,Q\in I_{\bar{\alpha}}\}$, hence by the set $\{g^{a_{i}-\alpha_{i}}\mid g\in N,i=1,\ldots,m\}$. The result below shows that the subgroup $(N)^{I_{\bar{\alpha}}}$ is definable in $G$ uniformly in $a_{1},\ldots,a_{m}$ and $(\alpha_{1},\ldots,\alpha_{m})$. Of course, integers are not elements of $G$, so instead of $(\alpha_{1},\ldots,\alpha_{m})$ we use their interpretations $a^{\alpha_{1}},\ldots,a^{\alpha_{m}}$ in ${\mathbb{Z}}_{a}$.
Lemma 4.7.
There is a formula $\phi(x,\bar{y},\bar{z})$ of the group language such that for any basis $(a_{1},\ldots,a_{m})$ of $A$ and any tuple $(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the formula
$\phi(x,a_{1},\ldots,a_{m},a^{\alpha_{1}},\ldots,a^{\alpha_{m}})$
defines in $G$ the subgroup $(N)^{I_{\bar{\alpha}}}$.
Proof.
Indeed, every element $u\in(N)^{I_{\bar{\alpha}}}$ can be written in the form
$u=g_{1}^{a_{1}-\alpha_{1}}\ldots g_{m}^{a_{m}-\alpha_{m}}=g_{1}^{a_{1}}g_{1}^{-\alpha_{1}}\ldots g_{m}^{a_{m}}g_{m}^{-\alpha_{m}},$
where $g_{1},\ldots,g_{m}\in N$. The elements $g_{i}^{\alpha_{i}}$ are the conjugations $g_{i}^{a_{i}}=a_{i}^{-1}g_{i}a_{i}$, so they are defined by formulas in $G$ with parameters $a_{i}$, $i=1,\ldots,m$. Exponents $g_{i}^{-\alpha_{i}}$ can be written by formulas in $G$ using the exponentiation formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})$ from Corollary 3.12.
Lemma 4.8.
Let $g,h\in N$ and $P\in{\mathbb{Z}}[\bar{a}]$. Then $g^{P}=h$ if and only if for every $\alpha_{1},\ldots\alpha_{m}\in{\mathbb{Z}}$ the following condition holds:
$g^{P(\alpha_{1},\ldots,\alpha_{m})}=h\ mod\ (N)^{I_{\bar{\alpha}}}.$ (15)
Proof.
Let $g^{p}=h$. Fix an arbitrary $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$. As we mentioned above
$P(a_{1},\ldots,a_{m})=P(\alpha_{1},\ldots,\alpha_{m})+U$
for some $U\in I_{\bar{\alpha}}$. Hence
$h=g^{P}=g^{p(\bar{\alpha})+U}=g^{P(\bar{\alpha})}g^{U}$
so the condition (15) holds.
Conversely, suppose for $g,h\in N$, $P\in{\mathbb{Z}}[\bar{a}]$, and arbitrary $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the condition (15) holds. We mentioned above, that $P=P(\bar{\alpha})\ mod\ I_{\bar{\alpha}}$, so (15) implies that
$g^{P}=hu$ (16)
for some $u\in(N)^{I_{\bar{\alpha}}}$. We need to show that $g^{P}=h$.
Decompose elements $g,h,u$ with respect to the basis $b_{1},\ldots,b_{n}$:
$g=\Pi_{i}b_{i}^{R_{i}},\ \ h=\Pi_{i}b_{i}^{S_{i}},\ \ u=\Pi_{i}b_{i}^{U_{i}},$
where $R_{i},S_{i}$ are Laurent polynomials from ${\mathbb{Z}}A$ and $U\in I_{\bar{\alpha}}$. Then the equality 16 implies
$\Pi_{i}b_{i}^{R_{i}P}=\Pi_{i}b_{i}^{S_{i}+U_{i}}$ (17)
Since $N$ is free with basis $b_{1},\ldots,b_{n}$ it follows from (17) that for $i=1,\ldots,n$
$R_{i}P=S_{i}+U_{i}$ (18)
in the ring ${\mathbb{Z}}A$. Let
$R_{i}=\frac{F_{i}}{K_{i}},\ \ S_{i}=\frac{H_{i}}{L_{i}},$
where $F_{i},H_{i}$ are ordinary polynomials from ${\mathbb{Z}}[\bar{a}]$, and $K_{i},L_{i}$ are non-zero monomials from ${\mathbb{Z}}[\bar{a}]$. Now, it follows from (18) that
$F_{i}PL_{i}=H_{i}K_{i}+U_{i}K_{i}L_{i}$
in the polynomial ring ${\mathbb{Z}}[\bar{a}]$.
Evaluating these polynomials at $a_{1}=\alpha_{1},\ldots,a_{m}=\alpha_{m}$ one gets
$F_{i}(\bar{\alpha})P(\bar{\alpha})L_{i}(\bar{\alpha})=H_{i}(\bar{\alpha})K_{i}(\bar{\alpha})$ (19)
since $U_{i}(\bar{\alpha})=0$. Hence by Lemma 4.6, we have equality in the polynomial ring ${\mathbb{Z}}[\bar{a}]$:
$F_{i}PL_{i}=H_{i}K_{i},$
which implies that
$\frac{F_{i}}{K_{i}}P=\frac{H_{i}}{L_{i}}$
in the ring of Laurent polynomials ${\mathbb{Z}}A$, that is
$R_{i}P=S_{i}$
for every $i=1,\ldots,n$. Hence $g^{P}=h$, as claimed. ∎
Now we can prove that in the group $G$ the action of ${\mathbb{Z}}A$ on $N$ is definable. Again, Laurent polynomials $Q\in{\mathbb{Z}}A$ are not elements of $G$, so we first represent them by their codes $k=\nu^{-1}(Q)$, and then represent the number $k$ as $a^{k}$ in the interpretation ${\mathbb{Z}}_{a}$ of $\mathbb{Z}$} in $G$. As above, we may choose $a=a_{1}$. The result is stated as follows.
Theorem 4.9.
There is a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ of group theory language such that for any $a\in A$, $g,h\in N$, $Q\in{\mathbb{Z}}A$, and $k=\nu^{-1}(Q)$ the following holds:
$g^{Q}=h\Longleftrightarrow G\models Act_{{\mathbb{Z}}A}(g,h,a^{k},a).$
Proof.
Fix $a\in A,a\neq 1$. We prove first, that the condition (15) from Lemma 4.8 can be written by formulas in $G$. Let $P\in{\mathbb{Z}}[a_{1},\ldots,a_{n}]$. Given the $k=\nu^{-1}{P}\in{\mathbb{Z}}$ and a tuple of numbers $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})$ one can compute the value $s=P(\bar{\alpha})\in{\mathbb{Z}}$, so there is a formula in arithmetic $Val_{\mathbb{Z}}(x_{1},x_{2},\bar{y})$ such that
$s=P(\bar{\alpha})\Longleftrightarrow{\mathbb{Z}}\models Val_{\mathbb{Z}}(\nu^{-1}(P),s,\bar{\alpha}).$
Hence, there is a formula $Val_{G}(x_{1},x_{2},\bar{y},z)$ of group theory such that
$a^{s}=a^{P(\bar{\alpha})}\Longleftrightarrow G\models Val_{G}(a^{\nu^{-1}(P)},a^{s},a^{\alpha_{1}},\ldots,a^{\alpha_{m}},a)$
To prove this one needs to rewrite the formula $Val_{\mathbb{Z}}(x_{1},x_{2},\bar{y})$ into a formula $Val_{G}(x_{1},x_{2},\bar{y},z)$ using the interpretation ${\mathbb{Z}}_{a}$ in $G$. The next step is to write a formula that for any $g,f\in N$, $P\in{\mathbb{Z}}[\bar{a}]$, and $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ states that $g^{P(\bar{\alpha})}=f$. This can be done using the formula $Exp_{G}(x_{1},x_{2},x_{3},y)$ from Theorem 4.5. Indeed, by Theorem 4.5
$g^{s}=f\Longleftrightarrow G\models Exp_{G}(g,f,a^{s},a).$
Now, given $a^{\nu^{-1}(P)}$, and $a^{s},a^{\alpha_{1}},\ldots,a^{\alpha_{m}},a$ the formula
$Val_{G}(a^{\nu^{-1}(P)},x_{2},a^{\alpha_{1}},\ldots,a^{\alpha_{m}},a)$
holds in $G$ only for $x_{2}=a^{p(\bar{\alpha})}$. Therefore,
$g^{P(\bar{\alpha})}=f\Longleftrightarrow G\models\exists x_{2}Val_{G}(a^{\nu^{-1}(P)},x_{2},a^{\alpha_{1}},\ldots,a^{\alpha_{m}},a)\wedge Exp_{G}(g,f,x_{2},a).$
We denote such a formula by $Exp_{N,{\mathbb{Z}}A}$.
Now we are able to write down by a formula the condition (15) from Lemma 4.8. Indeed, for any $\bar{\alpha}=(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the set $(N)^{I_{\bar{\alpha}}}$ is definable in $G$. The condition $g^{P(\bar{\alpha})}=f$ can be expressed by formula $Exp_{N,{\mathbb{Z}}A}$, so there is a formula $Act_{{\mathbb{Z}}[\bar{a}]}(x_{1},x_{2},x_{3},y_{1})$ of group language such that for any $a\in A$, $g,h\in N$, $P\in{\mathbb{Z}}[\bar{a}]$, and $k=\nu^{-1}(P)$ the following holds:
$g^{P}=h\Longleftrightarrow G\models Act_{{\mathbb{Z}}[\bar{a}]}(g,h,a^{k},a).$
This defines the action of the ring of ordinary polynomials ${\mathbb{Z}}[\bar{a}]$ on $N$.
The last step is to get a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ that defines the action of the ring of the Laurent polynomials ${\mathbb{Z}}A$ on $N$. Let $Q\in{\mathbb{Z}}A$. Then $Q=\frac{P}{M}$, where $P,M\in{\mathbb{Z}}[\bar{a}]$ and $M$ is a monomial. In this case, $g^{Q}=h$ if and only if $g^{P}=h^{M}$, and the exponentiation here by the polynomials $P$ and $M$, so the formula $Act_{{\mathbb{Z}}[\bar{a}]}(x_{1},x_{2},x_{3},y_{1})$ does it for us.
This proves the theorem.
Corollary 4.10.
There is a formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ of group language such that for any $u\in N$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
$u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}\Longleftrightarrow G\models\Lambda(u,c_{1}^{\nu(\bar{P})},\bar{c},\bar{u}),$ (20)
where $c_{1}^{\nu(\bar{P})}=(c_{1}^{\nu(P_{1})},\ldots,c_{1}^{\nu(P_{n})})$.
Proof.
Writing down the formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ suffices to define the powers $u_{j}^{P_{j}}$ by formulas in $G$. If $c_{1}=a_{1},\ldots,c_{m}=a_{m}$ then this is done in Theorem 4.9. Since for the top basis $\bar{c}$ there an automorphism $\phi\in Aut(G)$ which maps $\bar{a}$ to $\bar{c}$ component-wise, it follows that the result holds for $\bar{c}$ as well. ∎
Corollary 4.11.
There is a formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ of group language such that for any $g\in G$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{\gamma}=(\gamma_{1},\ldots,\gamma_{m})\in{\mathbb{Z}}^{m}$, and any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
$g=c_{1}^{\gamma_{1}}\ldots c_{m}^{\gamma_{m}}u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}\Longleftrightarrow G\models\Lambda_{G}(g,c_{1}^{\bar{\gamma}},c_{1}^{\nu(\bar{P})},\bar{c},\bar{u}),$ (21)
where $c_{1}^{\bar{\gamma}}=(c_{1}^{\gamma_{1}},\ldots,c_{1}^{\gamma_{m}})$ and $c_{1}^{\nu(\bar{P})}=(c_{1}^{\nu(P_{1})},\ldots,c_{1}^{\nu(P_{n})})$.
Proof.
The argument is similar to Corollary 4.10. The only difference is that to define the powers $c_{i}^{\gamma_{i}}$ by formulas in $G$ one uses Theorem 4.5. ∎
Theorem 4.12.
There is a formula $Base_{G}$ of group theory language (without parameters) which defines in $G$ the set of all bases of $G$.
Let $q=(c_{1},\ldots,c_{m},u_{1},\ldots,u_{n})$, where where $c_{i}\in G\smallsetminus N$ and $u_{i}\in N$. Denote $\bar{c}=(c_{1},\ldots,c_{m})$ and $\bar{u}=(u_{1},\ldots,u_{n})$. By Definition 2.5 $q$ is a basis of $G$ if and only if the following conditions hold:
  1. 1)
    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$,
  2. 2)
    $G=C\cdot N$,
  3. 3)
    NN is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
It suffices to describe a formula $Base_{G}(\bar{x},\bar{y})$ of the group theory language without parameters such that $G\models Base_{G}(\bar{c},\bar{u})$ if and only if $q$ satisfies the conditions 1)-3) above.
By Lemma 3.4 there is a formula $Base_{top}(x_{1},\ldots,x_{m})$ of group language (without parameters) such that for any $c_{1},\ldots,c_{m}\in G$ $G\models Base_{top}(c_{1},\ldots,c_{m})$ if and only if $(c_{1},\ldots,c_{m})$ is a top basis of $G$, in which case $c_{1},\ldots,c_{m}$ satisfy the conditions 1) and 2).
By Corollary 4.10 for $u\in N$ and $P_{1},\ldots,P_{n}\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ the condition $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ from (21) is described by the formula $\Lambda_{N}$. Using this one can easily write a formula that states that the element $u$ such that $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ is unique in $N$, which means that $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
This proves the theorem.
5 Bi-interpretability of $G$ and $\mathbb{Z}$}
Fix a tuple of parameters $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively. In this section we prove that $G$ and $\mathbb{Z}$} are bi-interpretable with parameters $\bar{p}$ in $G$.
Recall that $\mathbb{Z}$} is interpreted in $G$ on the cyclic subgroup $\langle a_{1}\rangle$, we denoted this interpretation ${\mathbb{Z}}_{a_{1}}=\Gamma(G,\bar{p})$ with the coordinate map $\mu_{\Gamma}:{\mathbb{Z}}_{a_{1}}\to{\mathbb{Z}}$ given by $a_{1}^{\gamma}\to\gamma$ for all $\gamma\in{\mathbb{Z}}$.
Meanwhile, $G$ is interpreted in $\mathbb{Z}$} by tuples of integers
$(\gamma_{1},\ldots,\gamma_{m},\nu(P_{1}),\ldots,\nu(P_{n})),$
where $\gamma_{i}\in{\mathbb{Z}}$, $P_{i}\in{\mathbb{Z}}[a_{1}^{\pm 1},\ldots,a_{m}^{\pm 1}]={\mathbb{Z}}A$, and $\nu(P_{i})$ is the code of the Laurent polynomial $P_{i}$. For simplicity we assume that the enumeration $\nu$ is such that the set $\nu({\mathbb{Z}}A)=\{\nu(P)\mid P\in{\mathbb{Z}}A\}$ is precisely $\mathbb{Z}$}. In this case the interpretation $G\simeq\Delta({\mathbb{Z}})$ of $G$ in $\mathbb{Z}$} is on the set ${\mathbb{Z}}^{m+n}$ with the coordinate map $\mu_{\Delta}:{\mathbb{Z}}^{m+n}\to G$ given by
$(\gamma_{1},\ldots,\gamma_{m},\nu(P_{1}),\ldots,\nu(P_{n}))\to g=a_{1}^{\gamma_{1}}\ldots a_{m}^{\gamma_{m}}b_{1}^{P_{1}}\ldots b_{n}^{P_{n}}.$
Theorem 5.1.
Let $G=B\,{\rm wr}\,A$.Then the following holds:
  1. 1)
    Let $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively. Then the group $G$ and the ring $\mathbb{Z}$} are bi-interpretable with each other through the interpretations $G\simeq\Delta({\mathbb{Z}})$ and ${\mathbb{Z}}\simeq\Gamma(G,\bar{p})$.
  2. 2)
    The group $G$ and the ring $\mathbb{Z}$} are regularly bi-interpretable with each other through the interpretations $G\simeq\Delta({\mathbb{Z}})$ and ${\mathbb{Z}}\simeq\Gamma(G,Base_{G})$, where $Base_{G}$ is the formula from Theorem 4.12 that defines the set of all bases in $G$.
Proof.
To prove 1) we need to show that the coordinate maps $\mu_{\Gamma\circ\Delta}$ and $\mu_{\Delta\circ\Gamma}$ are definable, correspondingly in $G$ (with parameters $\bar{p}$) and $\mathbb{Z}$}.
We consider $\mu_{\Gamma\circ\Delta}$, the other one is easy. Note that
$\mu_{\Gamma\circ\Delta}:(a_{1}^{\gamma_{1}},\ldots,a_{1}^{\gamma_{m}},a_{1}^{\nu(P_{1})},\ldots,a_{1}^{\nu(P_{n})})\to g=a_{1}^{\gamma_{1}}\ldots a_{m}^{\gamma_{m}}b_{1}^{P_{1}}\ldots b_{n}^{P_{n}}.$
By Corollary 4.11 the formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ does precisely what we need here, see (21).
2) follows from 1) by Theorem 4.12.
Theorem 5.1 implies Theorem 1.1 and Corollary 1.2 .
6 Condensed center-by-metabelian groups
6.1 The space of finitely generated marked groups.
We begin by recalling some definitions from [13]. For an abstract set $I$, we denote by $2^{I}$ the space of all subsets of $I$ endowed with the product topology (equivalently, the topology of pointwise convergence of the indicator functions). If $G$ is a group, we denote by $\mathcal{N}(G)$ the set of all normal subgroups of $G$ endowed with the subspace topology induced by the inclusion $\mathcal{N}(G)\subseteq 2^{G}$. Thus, the base of neighborhoods of $N\in\mathcal{N}(G)$ consists of the sets
$U(\mathcal{F},N)=\{M\leq G\mid M\cap\mathcal{F}=N\cap\mathcal{F}\},$ (22)
where $\mathcal{F}$ ranges in the set of all finite subsets of $G$.
Let $n\in\mathbb{N}$. An $n$-generated marked group (or simply a marked group) is a pair $(G,A)$, where $G$ is a group and $(a_{1},\ldots,a_{n})\subseteq G^{n}$ is an $n$-tuple such that $G$ is generated by $\{a_{1},\ldots,a_{n}\}$. Two $n$-generated marked groups $(G,(a_{1},\ldots,a_{n}))$ and $(H,(b_{1},\ldots,b_{n}))$ are isomorphic if the map sending $a_{i}$ to $b_{i}$ for all $i=1,\ldots,n$ extends to an isomorphism $G\to H$. The set of all isomorphism classes of $n$-generated marked groups is denoted by $\mathcal{G}_{n}$. Following the common practice, we abuse the notation and terminology by not distinguishing between marked groups and their isomorphism classes.
For every $n\in\mathbb{N}$, let $F_{n}$ be the free group of rank $n$ with a fixed basis $X=\{x_{1},\ldots,x_{n}\}$. For every $(G,A)\in\mathcal{G}_{n}$, where $A=(a_{1},\ldots,a_{n})$, let $\varepsilon_{(G,A)}\colon F_{n}\to G$ be the homomorphism such that
$\varepsilon_{(G,A)}(x_{i})=a_{i}\;\;\;\forall\,i=1,\ldots,n.$
The map $(G,A)\to\operatorname{Ker}\varepsilon_{(G,A)}$ is a well-defined bijection between $\mathcal{G}_{n}$ and $\mathcal{N}(F_{n})$. The topology on $\mathcal{G}_{n}$ is defined as the pull-back topology under this map; that is, a sequence $\{(G_{i},A_{i})\}_{i\in\mathbb{N}}$ converges to $(G,A)$ in $\mathcal{G}_{n}$ if and only if the sequence $\{\operatorname{Ker}\varepsilon_{(G_{i},A_{i})}\}_{i\in\mathbb{N}}$ converges to $\operatorname{Ker}\varepsilon_{(G,A)}$ in $\mathcal{N}(F_{n})$.
Example 6.1.
We have $\lim\limits_{n\to\infty}(\mathbb{Z}_{n},\{1\})=(\mathbb{Z},\{1\})$ in $\mathcal{G}_{1}$.
It is easy to see that the map $(G,(a_{1},\ldots,a_{n}))\mapsto(G,(a_{1},\ldots,a_{n},1))$ defines a continuous embedding $\mathcal{G}_{n}\to\mathcal{G}_{n+1}$.
Definition 6.2.
The topological union
$\mathcal{G}=\bigcup_{n=1}^{\infty}\mathcal{G}_{n}$
is called the space of finitely generated marked groups.
By the definition of a topological union, a set $U\subseteq\mathcal{G}$ is open if and only if $U\cap\mathcal{G}_{n}$ is open in $\mathcal{G}_{n}$ for all $n\in\mathbb{N}$. Alternatively, the topology on $\mathcal{G}$ can be as follows: a sequence $\{(G_{i},A_{i})\}_{i\in\mathbb{N}}$ converges to a marked group $(G,A)\in\mathcal{G}_{n}$ if and only if there exists $M\in\mathbb{N}$ such that $(G_{i},A_{i})\in\mathcal{G}_{n}$ for all $i\geq M$, and the subsequence $\{(G_{i},A_{i})\}_{i=M}^{\infty}$ converges to $(G,A)$ in $\mathcal{G}_{n}$.
Recall that a topological space is Polish if it is separable and completely metrizable. For the result below, see [13] and [34, Proposition 3.7].
Proposition 6.3.
The space $\mathcal{G}$ is Polish.
We will need the following elementary yet useful observation.
Lemma 6.4 ([28, Lemma 2.2]).
Let $G$ be a group generated by a finite set $X$. Given a normal subgroup $N\lhd G$, we let $X_{N}$ denote the image of $X$ under the natural homomorphism $G\to G/N$. The map $\mathcal{N}(G)\to\mathcal{G}$ defined by the formula $N\mapsto(G/N,X_{N})$ for all $N\in\mathcal{N}(G)$ is continuous.
6.2 The main construction
Throughout this section, we use the standard notation $x^{y}=y^{-1}xy$ and $x,y]=x^{-1}y^{-1}xy$. We begin by constructing an auxiliary central extension of $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$, which is reminiscent of the center-by-metabilian groups discussed in [15] (see also [11]).
Let
$A=\{a_{i},b_{i},c_{i}\mid i\in\mathbb{Z}\}$ (23)
and
[…]
where the relations are imposed for all $\forall\,i,j\in\mathbb{Z}$.
Lemma 6.5.
Let $G$ be the group defined above.
  1. (a)
    The center $Z(G)$ is a free group in the variety of abelian groups of exponent $2$, and $\{c_{k}\mid k\in\mathbb{Z}\}$ is a basis in $Z(G)$.
  2. (b)
    $G/Z(G)\cong\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$.
Proof.
Let $N$ be the free nilpotent group of rank $2$ with basis $\{\tilde{a}_{i},\tilde{b}_{i}\mid i\in\mathbb{Z}\}$. It is well-known that $[N,N]=Z(N)$ is a free abelian group freely generated by the basic commutators $\tilde{a}_{i},\tilde{a}_{j}]$, $\tilde{b}_{i},\tilde{b}_{j}]$ for $j<j$, $i,j\in\mathbb{Z}$, and $\tilde{a}_{i},\tilde{b}_{j}]$ for $i,j\in\mathbb{Z}$ (see [29, Ch. 3, Sec. 1] and references therein). Let $M\leq[N,N]$ be the subgroup generated by the elements
$\tilde{a}_{i},\tilde{a}_{j}],\;\;[\tilde{b}_{i},\tilde{b}_{j}],\;\;[\tilde{a}_{i},\tilde{b}_{j}]^{2},\;\;[\tilde{a}_{i},\tilde{b}_{j}][\tilde{a}_{i+\ell},\tilde{b}_{j+\ell}]$
for all $i,j,\ell\in\mathbb{Z}$. It is straightforward to verify that $[N,N]/M$ is a free group in the variety of abelian groups of exponent $2$, freely generated by elements $c_{k}=[\tilde{a}_{0},\tilde{b}_{k}]M$, $k\in\mathbb{Z}$.
Since $M$ is central in $N$, we can consider the quotient group $N/M$. We denote the images of $\tilde{a}_{i}$ and $\tilde{b}_{i}$ ($i,j\in\mathbb{Z}$) in $N/M$ by $a_{i}$ and $b_{j}$, respectively. Our construction yields the presentation
[…]
where $A$ is given by (23) and the indices $i$, $j$ independently range in $\mathbb{Z}$.
The map $\tilde{a}_{i}\mapsto\tilde{a}_{i+1}$, $\tilde{b}_{i}\mapsto\tilde{b}_{i+1}$ ($i\in\mathbb{Z}$) extends to an automorphism $N\to N$, which fixes $M$ setwise. Therefore, it induces an automorphism $\phi\in\operatorname{Aut}(N/M)$ such that $\phi(a_{i})=a_{i+1}$ and $\phi(b_{i})=b_{i+1}$ for all $i\in\mathbb{Z}$. Note also that for every $i\in\mathbb{Z}$, we have
$\phi(c_{i})=\phi([\tilde{a}_{0},\tilde{b}_{i}]M)=[\tilde{a}_{1},\tilde{b}_{i+1}]M=[\tilde{a}_{0},\tilde{b}_{i}]M=c_{i}.$
Hence, the group $G$ given by (24) is naturally isomorphic to the mapping torus $N/M\rtimes_{\phi}\mathbb{Z}$. In particular, $N/M$ naturally embeds in $G$.
The subgroup $N,N]/M=\langle\{c_{i}\mid i\in\mathbb{Z}\}\rangle\leq G$ is central in $G$ as all generators of $G$ commute with $c_{i}$ for all $i\in\mathbb{Z}$. The quotient group $G/([N,N]/M)$ has the presentation
$\langle t,a_{i},b_{i}\mid[a_{i},a_{j}]=[b_{i},b_{j}]=[a_{i},b_{j}]=1,\;a_{i}^{t}=a_{i+1},\;b_{i}^{t}=b_{i+1}\rangle,$
which is one of the standard presentations of the wreath product $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$. Since the latter group is centerless, we have $Z(G)=[N,N]/M$, and conditions (a) and (b) follow. ∎
For every subset $S\subseteq\mathbb{Z}$, we define a subgroup $Z_{S}\leq Z(G)$ by the formula
$Z_{S}=\Big\langle\{c_{i}\mid i\in\mathbb{Z}\setminus S\}\cup\{c_{j}c_{k},\mid,\;j,k\in S\}\Big\rangle$
and let
$G_{S}=G/Z_{S}.$
For instance, we have $G_{\emptyset}\cong\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$.
Lemma 6.6.
For every non-empty subset $S\subseteq\mathbb{Z}$, the group $G_{S}$ is a central extension of the form
$1\longrightarrow\mathbb{Z}_{2}\longrightarrow G_{S}\longrightarrow\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}\longrightarrow 1,$
where the kernel of the extension is generated by $c_{i}$ for any $i\in S$.
Proof.
Part (a) of Lemma 6.5 and the definition of $Z_{S}$ imply that the image of $Z(G)$ in $G_{S}$ is isomorphic to $Z(G)/Z_{S}\cong\mathbb{Z}_{2}$. The result now follows from part (b) of Lemma 6.5. ∎
We consider the shift action of $\mathbb{Z}$ on $2^{\mathbb{Z}}$ by homeomorphisms defined by
$n\circ S=\{s-n\mid s\in S\}$
for all $n\in\mathbb{Z}$ and all $S\in 2^{\mathbb{Z}}$.
Lemma 6.7.
For every $S\subseteq\mathbb{Z}$ and every $n\in\mathbb{Z}$, the group $G_{S}$ is generated by the elements $a=a_{0}$, $b_{n}$, and $t$; in these generators, it has the presentation
$G_{S}=\left\langle a,b_{n},t\;\left|\;\begin{array}[]{c}{[a,a^{t^{i}}]=[b_{n},b_{n}^{t^{i}}]=1\;\;\forall\,i\in\mathbb{Z}}\\ {[a,[a,b_{n}^{t^{i}}]]=[b_{n},[a,b_{n}^{t^{i}}]]=[t,[a,b_{n}^{t^{i}}]]=[a,b_{n}^{t^{i}}]^{2}=1\;\;\forall\,i\in\mathbb{Z}}\\ {[a,b_{n}^{t^{j}}]=1\;\;\forall\,j\in n\circ(\mathbb{Z}\setminus S),\;\;[a,b_{n}^{t^{k}}]=[a,b_{n}^{t^{\ell}}]\;\;\forall\,k,\ell\in n\circ S}\end{array}\right.\right\rangle$. (25)
Proof.
To obtain the required presentation of $G_{S}$, we proceed as follows. We first note that it suffices to impose the relations in the second and third rows of (24) for $i=0$ only; all other relations are redundant and we omit them from the presentation of $G$. Further, we obtain a presentation of $G_{S}$ by adding the relations $c_{i}=1$ and $c_{j}=c_{k}$ for all $i\in\mathbb{Z}\setminus S$ and $j,k\in S$. We rename $a_{0}$ by $a$ and exclude all redundant generators using the relations $a_{i}=a^{t^{i}}$, $b_{i}=b_{n}^{t^{i-n}}$, and $c_{i}=[a,b_{n}^{t^{i-n}}]$. In particular, relations of the form $c_{i}=1$ and $c_{j}=c_{k}$ become $a,b_{n}^{t^{i-n}}]=1$ and $a,b_{n}^{t^{j-n}}]=[a,b_{n}^{t^{k-n}}]$, respectively, where $i\in\mathbb{Z}\setminus S$ and $j,k\in S$. We replace the exponent ini-n with $i\in n\circ(\mathbb{Z}\setminus S)$ and knk-n with $k\in n\circ S$. Finally, removing the redundant relations occurring from the first row of relations in (24), we obtain (25). ∎
Corollary 6.8.
For every $S\subseteq\mathbb{Z}$ and every $n\in\mathbb{Z}$, we have $G_{S}\cong G_{n\circ S}$.
Proof.
By Lemma 6.7, the required isomorphism is given by the map $a\mapsto a$, $b_{0}\mapsto b_{n}$, $t\mapsto t$. ∎
Lemma 6.9.
The map $f\colon 2^{\mathbb{Z}}\to\mathcal{G}$ given by $f(S)=(G_{S},\{a,b_{0},t\})$ is injective and continuous.
Proof.
To prove injectivity, let $S$ and $T$ be two distinct subsets of $\mathbb{Z}$. Without loss of generality, we can assume that there exists $i\in S\setminus T$. By Lemma 6.7, the relation $a,b_{0}^{t^{i}}]=1$ holds in $G_{T}$. On the other hand, the element $a,b_{0}^{t^{i}}]=c_{i}$ generates $Z(G_{S})\cong\mathbb{Z}_{2}$ by Lemma 6.6. In particular, $a,b_{0}^{t^{i}}]\neq 1$ in $G_{S}$. By the definition of marked groups, this implies that $(G_{S},\{a,b_{0},t\})$ and $(G_{T},\{a,b_{0},t\})$ represent distinct elements of $\mathcal{G}$.
We now show that the map $f$ is continuous. To this end, we represent it as the composition of maps
$2^{\mathbb{Z}}\stackrel{{\scriptstyle\alpha}}{{\to}}\mathcal{N}(G)\stackrel{{\scriptstyle\beta}}{{\to}}\mathcal{G},$
where $\alpha(S)=Z_{S}\lhd G$ for all $S\subseteq\mathbb{Z}$, and $\beta(N)=(G/N,\{a_{0}N,b_{0}N,tN\})$ for all $N\lhd G$. Observe that $G$ is generated by $a_{0}$, $b_{0}$, and $t$; hence, the map $\beta$ is well-defined and continuous by Lemma 6.4.
It remains to show that the map $\alpha$ is continuous. By the definition of the topologies on $2^{\mathbb{Z}}$ and $\mathcal{N}(G)$, this amounts to showing that for every finite subset $\mathcal{F}\subseteq G$, there exists a finite subset $\mathcal{E}\subseteq\mathbb{Z}$ such that for every $S,T\in 2^{\mathbb{Z}}$ satisfying
$S\cap\mathcal{E}=T\cap\mathcal{E},$ (26)
we have $Z_{S}\cap\mathcal{F}=Z_{T}\cap\mathcal{F}$. We fix a finite subset $\mathcal{F}\subseteq G$. By part (a) of Lemma 6.5, every element $z\in\mathcal{F}\cap Z(G)$ can be uniquely decomposed as
$z=\prod_{i\in\mathcal{K}(z)}c_{i}$
for some finite subset $\mathcal{K}(z)\subseteq\mathbb{Z}$. We let
$\mathcal{E}=\bigcup_{z\in\mathcal{F}\cap Z(G)}\mathcal{K}(z).$
Suppose that $S$ and $T$ are subsets of $\mathbb{Z}$ satisfying (26). By the definition of $Z_{S}$, every element $z\in Z_{S}$ can be decomposed as
$z=\left(\prod_{i\in\mathcal{P}}c_{i}\right)\cdot\left(\prod_{(j,k)\in\mathcal{R}}c_{j}c_{k}\right)$ (27)
for some finite subsets $\mathcal{P}\subseteq\mathbb{Z}\setminus S$ and $\mathcal{R}\subseteq S\times S$. We take such a decomposition with the minimal possible number of factors, which ensures that every $c_{i}$ appears in the decomposition at most once (for example, if the second product contains $c_{j}c_{k}$ and $c_{j}c_{\ell}$ for some $j,k,l\in S$, we can use the equality $c_{j}c_{k}c_{j}c_{\ell}=c_{k}c_{\ell}$ to reduce the number of factors). Since $\{c_{i}\mid i\in\mathbb{Z}\}$ is a basis in $Z(G)$, an element $c_{i}$ occurs in (27) if and only if $i\in\mathcal{K}(z)\subseteq\mathcal{E}$. Combining this with (26), we obtain that $\mathcal{P}\subseteq\mathbb{Z}\setminus T$ and $\mathcal{R}\subseteq T\times T$. Consequently, $z\in Z_{T}$. This proves the inclusion $Z_{S}\cap\mathcal{F}\leq Z_{T}\cap\mathcal{F}$. The proof of the opposite inclusion is symmetric, so we obtain $Z_{S}\cap\mathcal{F}=Z_{T}\cap\mathcal{F}$, as required. This concludes the proof of the continuity of $\alpha$ and the lemma. ∎
Recall that the action of a group $H$ by homeomorphisms on a topological space $X$ is said to be topologically transitive if, for any non-empty open subsets $U,V\subseteq X$, there exists $g\in G$ such that $gU\cap V\neq\emptyset$. If $X$ is Polish, this condition is equivalent to the existence of a dense orbit.
Proof of Theorem 1.5.
It is well-known and straightforward to prove that the shift action of $\mathbb{Z}$ on $2^{\mathbb{Z}}$ is topologically transitive. Thus, there exists $S\in 2^{\mathbb{Z}}$ with a dense orbit. In particular, the orbit of $S$ has no isolated points. By Lemma 6.9 and Corollary 6.8, the image of the orbit of $S$ under the map $f$ has no isolated points and belongs to the isomorphism class $G_{S}]\subseteq\mathcal{G}$. It follows that $G_{S}]$ is non-discrete. Since the isomorphism class of every finitely generated group in $\mathcal{G}$ is either discrete or has no isolated points (see [34, Corollary 6.1]), we obtain that $G_{S}$ is condensed. ∎
2.3. Elementarily equivalent center-by-metabelian groups.
We are now ready to prove the corollaries announced in the introduction.
Proof of Corollary 1.6.
This is an immediate consequence of Theorem 1.5 and [34, Proposition 2.7] applied to the subset of $\mathcal{G}$ consisting of all finitely generated, center-by-metabelian marked groups. Indeed, the aforementioned proposition states that an isomorphism-invariant closed subset $S$ of $\mathcal{G}$ contains a condensed group if and only if the isomorphism relation on $S$ is not smooth. ∎
Proof of Corollary 1.7.
As observed in [34, Proposition 6.2], finitely generated abelian-by-nilpotent (in particular, finitely generated metabelian) groups cannot be condensed. Thus, $\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$ is not condensed. Combining this with Theorem 1.5, we obtain the desired result. ∎
The proof of Corollary 1.8 follows the same idea as that of Proposition 2.3 in [34] and involves proving a zero-one law for the set $f(2^{\mathbb{Z}})=\{(G_{S},\{a,b_{0},t\})\mid S\subseteq\mathbb{Z}\}$ analogous to Theorem 2.2 in [34]. Unfortunately, we cannot apply the latter theorem in our settings as $f(2^{\mathbb{Z}})$ is not isomorphism-invariant in the terminology of [34]. To address this problem, we first prove the proposition below, which seems to be of independent interest.
Definition 6.10.
Let $H$ be a group acting on the space $\mathcal{G}$. We say that the action is isomorphism-preserving if, for every marked group $(K,X)\in\mathcal{G}$, the $H$-orbit of $(K,X)$ belongs to the isomorphism class $[K]$.
Recall that a subset of a topological space is perfect if it is closed and has no isolated points.
Proposition 6.11.
Let $\mathcal{C}$ be a perfect subspace of $\mathcal{G}$. Suppose that $Homeo(\mathcal{C})$ contains an isomorphism-preserving, topologically transitive subgroup. Then $\mathcal{C}$ contains a subset $\mathcal{C}_{0}$ of cardinality $|\mathcal{C}_{0}|=2^{\aleph_{0}}$ such that, for any $(K,X),(L,Y)\in\mathcal{C}_{0}$, we have $K\equiv L$.
Proof.
The proof of the proposition repeats the main steps of the proofs of Theorem 2.2 and Proposition 2.3 in [34], so we keep it brief.
For every first-order sentence $\sigma$ in the language of groups $\mathcal{L}$, let
${\rm Mod}_{\mathcal{C}}(\sigma)=\{(K,X)\in\mathcal{C}\mid K\models\sigma\}.$
By [34, Proposition 5.1], ${\rm Mod}_{\mathcal{C}}(\sigma)$ is always a Borel subset of $\mathcal{C}$. Since $\mathcal{C}$ is a closed subspace of the Polish space $\mathcal{G}$, it is a Polish space itself. Let $H\leq Homeo(\mathcal{S})$ by an isomorphism-preserving, topologically transitive subgroup. By the zero-one law for topologically transitive group actions on Polish spaces (see [18, Theorem 8.46]), every $H$-invariant Borel subset of $\mathcal{C}$ is either meager or comeager. In particular, ${\rm Mod}_{\mathcal{C}}(\sigma)$ is either meager or comeager for every sentence $\sigma$ in $\mathcal{L}$.
Let $Th^{gen}(\mathcal{C})$ denote the set of all sentences $\sigma$ in $\mathcal{L}$ such that ${\rm Mod}_{\mathcal{C}}(\sigma)$ is comeager in $\mathcal{C}$. Since the intersection of a countable collection of comeager sets is comeager, there is a comeager set $\mathcal{C}_{0}\subseteq\mathcal{C}$ such that, for every $(K,X)\in\mathcal{C}_{0}$, the group $K$ is a model of $Th^{gen}(\mathcal{C})$. On the other hand, for every sentence $\sigma$, we have ${\rm Mod}_{\mathcal{C}}(\sigma)\cup{\rm Mod}_{\mathcal{C}}(\lnot\sigma)=\mathcal{C}$. By the Baire category theorem, $\mathcal{C}$ cannot be represented as a usion of two meager sets. Hence, either ${\rm Mod}_{\mathcal{C}}(\sigma)$ or ${\rm Mod}_{\mathcal{C}}(\lnot\sigma)$ must be comeager. This implies that the theory $Th^{gen}(\mathcal{C})$ is complete. Since all models of a complete theory are elementarily equivalent, we have $K\equiv L$ for all $(K,X),(L,Y)\in\mathcal{C}_{0}$.
Finally, recall that every comeager subset of a Polish space contains a dense $G_{\delta}$-subset by the Baire category theorem. In particular, $\mathcal{C}_{0}$ contains a dense $G_{\delta}$-subset $\mathcal{C}_{1}$. Since $\mathcal{C}_{0}$ is perfect, $\mathcal{C}_{1}$ has no isolated points. Every $G_{\delta}$-subspace of a Polish space is Polish (see [18, Theorem I.3.11]), hence $\mathcal{C}_{1}$ is Polish. Being a Polish space without isolated points, $\mathcal{C}_{1}$ has the cardinality of the continuum by [18, Corollary I.6.3]. Therefore, we have $|\mathcal{C}_{0}|=2^{\aleph_{0}}$. ∎
Proof of Corollary 1.8.
Let $\mathcal{C}=f(2^{\mathbb{Z}})$, where $f$ is the map from Lemma 6.9. Note that $\mathcal{C}$ consists of groups of the form (1) by Lemma 6.6. Since $2^{\mathbb{Z}}$ is compact and $f$ is continuous, $\mathcal{C}$ is compact; in particular, $\mathcal{C}$ is a closed subset of $\mathcal{G}$. Note also that $\mathcal{C}$ has no isolated points being the image of a space without isolated points under a continuous and injective map.
The shift action of $\mathbb{Z}$ on $2^{\mathbb{Z}}$ induces an action of $\mathbb{Z}$ on $\mathcal{C}$ by the rule
$n(G_{S},\{a,b_{0},t\})=(G_{n\circ S},\{a,b_{0},t\}).$
Since the map $f\colon 2^{\mathbb{Z}}\to\mathcal{C}$ is bijective and continuous with compact domain, it is a homeomorphism; this implies that the action of $\mathbb{Z}$ on $\mathcal{C}$ defined above is continuous; in addition it is topologically transitive as so is the action of $\mathbb{Z}$ on $2^{\mathbb{Z}}$. Further, the action of $\mathbb{Z}$ on $\mathcal{C}$ is isomorphism-preserving by Corollary 6.8.
Applying Proposition 6.11, we obtain a subset $\mathcal{C}_{0}$ of cardinality $|\mathcal{C}_{0}|=2^{\aleph_{0}}$ such that, for any $(G_{S},\{a,b_{0},t\}),(G_{T},\{a,b_{0},t\})\in\mathcal{C}_{0}$, we have $G_{S}\equiv G_{T}$. Clearly, for every finitely generated group $K$, the isomorphism class $[K]$ in $\mathcal{G}$ is countable. Therefore, there are $2^{\aleph_{0}}$ pairwise non-isomorphic, elementarily equivalent groups among $G_{S}$, $S\subseteq\mathbb{Z}$. ∎

### 18 Mar 2026 v2

@@
 In this paper we prove two main results. The first is the following theorem, which answers Problem 4.3 from [22] in the affirmative.
-Theorem 1.1.
-Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is regularly injectively bi-interpretable with $\mathbb{Z}$.
+Theorem $1.$
+Let AA and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is regularly injectively bi-interpretable with $\mathbb{Z}$.
 The proof of Theorem 1.1 is divided into four parts: interpretability of $G$ in $\mathbb{Z}$} in Section 3.1, interpretability of $\mathbb{Z}$} in $G$ in Sectiom 3.2, bi-interpretability of $G$ and $\mathbb{Z}$} with parameters, and finally, regular bi-interpretability of $G$ and $\mathbb{Z}$} in Sections 4, 5.
 Corollary 1.2.
-Let $A$ and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is QFA.
+Let AA and $B$ be free abelian groups of finite ranks. Then $B\,{\rm wr}\,A$ is QFA.
 Theorem 1.1 and results from [8], Section 5.8 imply the following.
@@
 To state the second theorem, we need some notation and terminology.
-Recall that the space of finitely generated marked groups $\mathcal{G}$ can be informally defined as the set of all pairs $(G,A)$, where $G$ is a group and $A$ is an ordered finite generating set of $G$, endowed with the topology induced by local convergence of Cayley graphs.
+Recall that the space of finitely generated marked groups $\mathcal{G}$ can be informally defined as the set of all pairs $(G,A)$, where $G$ is a group and AA is an ordered finite generating set of $G$, endowed with the topology induced by local convergence of Cayley graphs.
 Given a finitely generated group $G$, we denote by $G]\subseteq\mathcal{G}$ its isomorphism class; that is,
@@
 The first one concerns the complexity of the isomorphism relation.
-Recall that an equivalence relation $E$ on a topological space $X$ is called smooth if there is a Polish space $P$ and a Borel map $\beta\colon X\to P$ such that for any $x,y\in X$, we have $xEy$ if and only if $\beta(x)=\beta(y)$.
-In [46], Williams proved that the isomorphism relation on the space of $3$-step solvable, finitely generated, marked groups is non-smooth. We strengthen this result as follows.
+Recall that an equivalence relation $E$ on a topological space $X$ is called smooth if there is a Polish space $P$ and a Borel map $\beta\colon X\to P$ such that for any $x,y\in X$, we have xEyxEy if and only if $\beta(x)=\beta(y)$.
+In [46], Williams proved that the isomorphism relation on the space of 33-step solvable, finitely generated, marked groups is non-smooth. We strengthen this result as follows.
 Corollary 1.6.
@@
 $A=\{(b_{1},\ldots,b_{n})\in B^{n}\mid\mathbb{B}\models\phi(b_{1},\ldots,b_{n})\}.$
-In this case we denote $A$ by $\phi(B^{n})$ or $\phi({\mathbb{B}})$ and say that $\phi$ defines $A$ in $\mathbb{B}$.
-Similarly, an operation $f$ on the subset $A$ is definable in $\mathbb{B}$ if its graph is definable in $\mathbb{B}$.
+In this case we denote AA by $\phi(B^{n})$ or $\phi({\mathbb{B}})$ and say that $\phi$ defines AA in $\mathbb{B}$.
+Similarly, an operation $f$ on the subset AA is definable in $\mathbb{B}$ if its graph is definable in $\mathbb{B}$.
 A constant $c$ is definable if the relation $x=c$ is definable.
@@
 In the notation above if $\sim$ is a definable equivalence relation on a definable subset $A\subseteq B^{n}$ then we say that the quotient set $A/\sim$ is interpretable in $\mathbb{B}$.
-Furthermore, an operation $f$ or a predicate $P$ on the quotient set $A/\sim$ is interpretable in $\mathbb{B}$ if the full preimage of its graph in $A$ is definable in $\mathbb{B}$.
-For example, if $N$ is a normal definable subgroup of a group $G$, then the equivalence relation $x\sim y$ on $G$ given by $xN=yN$ is definable in $G$, so the quotient set $G/N$ of all right cosets of $N$ is interpretable in $G$.
+Furthermore, an operation $f$ or a predicate $P$ on the quotient set $A/\sim$ is interpretable in $\mathbb{B}$ if the full preimage of its graph in AA is definable in $\mathbb{B}$.
+For example, if NN is a normal definable subgroup of a group $G$, then the equivalence relation $x\sim y$ on $G$ given by $xN=yN$ is definable in $G$, so the quotient set $G/N$ of all right cosets of NN is interpretable in $G$.
 It is easy to see that the multiplication induced from $G$ on $G/N$ is also interpretable in $G$.
@@
 Now we discuss a very strong version of mutual interpretability of two structures, so-called bi-interpretability.
-Definition 2.2.
+Definition $2.$
 Two algebraic structures $\mathbb{A}$} and $\mathbb{B}$} are called bi-interpretable (with parameters) in each other if the following conditions hold:
@@
 Put ${\mathbb{N}}_{0}=\lambda({\mathbb{N}})$.
-Then the restriction $\nu_{0}$ of $\nu$ to ${\mathbb{N}}_{0}$ gives a bijective enumeration $\nu_{0}:{\mathbb{N}}_{0}\to A$ such that the full $\nu_{0}$-preimages of the basic operations and predicates on $A$ are computable in $\mathbb{N}$}.
+Then the restriction $\nu_{0}$ of $\nu$ to ${\mathbb{N}}_{0}$ gives a bijective enumeration $\nu_{0}:{\mathbb{N}}_{0}\to A$ such that the full $\nu_{0}$-preimages of the basic operations and predicates on AA are computable in $\mathbb{N}$}.
 Since $\nu_{0}$ is a bijection, this construction gives an $L$-structure denoted by ${\mathcal{A}}_{\nu}$ on ${\mathbb{N}}_{0}$ that is isomorphic to $\mathcal{A}$} with respect to $\nu_{0}$.
@@
 2.3 Wreath products of free abelian groups
-Let $A$ and $B$ be free abelian groups with bases $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$, respectively.
-If not stated otherwise, we view the groups $A$ and $B$ in multiplicative notation.
-Denote by $G=B\,{\rm wr}\,A$ the restricted wreath product of $A$ and $B$.
+Let AA and $B$ be free abelian groups with bases $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$, respectively.
+If not stated otherwise, we view the groups AA and $B$ in multiplicative notation.
+Denote by $G=B\,{\rm wr}\,A$ the restricted wreath product of AA and $B$.
 We fix this notation throughout the entire paper.
-By construction, the group $G$ is a semidirect product $N\rtimes A$, where $N$ is the direct product $N\simeq\Pi_{a\in A}B_{a}$ of isomorphic copies $B_{a}$ of $B$, termed the base group of this wreath product, and $A$, termed the active group, acts on $N$ by shifting the copies $B_{a}$ in $N$.
+By construction, the group $G$ is a semidirect product $N\rtimes A$, where NN is the direct product $N\simeq\Pi_{a\in A}B_{a}$ of isomorphic copies $B_{a}$ of $B$, termed the base group of this wreath product, and AA, termed the active group, acts on NN by shifting the copies $B_{a}$ in NN.
 Namely if $f\in\Pi_{a\in A}B_{a}$ is viewed as a function $f:A\to\cup_{a\in A}B_{a}$, then the action of $a\in A$ on $f$ results in a function $f^{a}(x)=f(ax)$ for any $x\in A$.
@@
 Note that in this case $f^{a}=a^{-1}fa$ - the conjugation of $f$ by $a$.
-The groups $A$ and $B$ canonically embed into $G$ via the maps $a\in A\to a\cdot 1_{N}\in G$ and $b\in B\to f_{b}$, where $f_{b}(x)=1$ for $x\neq 1$ and $f_{b}(1)=b$.
-We often identify $A$ and $B$ with their images in $G$, so $B$ is identified with the copy $B_{1}$.
+The groups AA and $B$ canonically embed into $G$ via the maps $a\in A\to a\cdot 1_{N}\in G$ and $b\in B\to f_{b}$, where $f_{b}(x)=1$ for $x\neq 1$ and $f_{b}(1)=b$.
+We often identify AA and $B$ with their images in $G$, so $B$ is identified with the copy $B_{1}$.
 We list below some known properties of wreath products $G$.
   1. 1)
-    The subgroup $N$ of $G$ is precisely the normal closure $N=ncl(B)$ of $B$ in $G$.
+    The subgroup NN of $G$ is precisely the normal closure $N=ncl(B)$ of $B$ in $G$.
   2. 2)
-    The action of $A$ on $N$ extends linearly to the action of the integer group ring ${\mathbb{Z}}A$ on $N$, so for $u=\Sigma n_{i}a_{i}\in{\mathbb{Z}}A$ and $t\in N$ $t^{u}=\Pi(t^{n_{i}})^{a_{i}}$.
-This makes $N$ into ${\mathbb{Z}}A$-module.
-Observe, that we use here the multiplicative notation for the action of ${\mathbb{Z}}A$ on $N$, which matches naturally with the conjugation notation in $G$.
+    The action of AA on NN extends linearly to the action of the integer group ring ${\mathbb{Z}}A$ on NN, so for $u=\Sigma n_{i}a_{i}\in{\mathbb{Z}}A$ and $t\in N$ $t^{u}=\Pi(t^{n_{i}})^{a_{i}}$.
+This makes NN into ${\mathbb{Z}}A$-module.
+Observe, that we use here the multiplicative notation for the action of ${\mathbb{Z}}A$ on NN, which matches naturally with the conjugation notation in $G$.
   3. 3)
@@
 For convenience, in this and similar cases, we write sometimes $p=\bar{a}\cdot\bar{b}$, where $\bar{a}=(a_{1},\ldots,a_{m})$, $\bar{b}=(b_{1},\ldots,b_{n})$ and $\cdot$ is concatenation of tuples.
-Respectively, we write $s_{p}(g)=\bar{\gamma}\cdot\bar{P}$ for the tuple of coordinates of $g$ relative to $p$.
+Respectively, we write $s_{p}(g)=\bar{\gamma}\cdot\bar{P}$ for the tuple of coordinates of $g$ relative to pp.
 Using this notation we sometimes write $g={\bar{a}}^{\bar{\gamma}}\cdot{\bar{b}}^{\bar{P}}$.
@@
   2. 2)
-    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$, $G=C\cdot N$, $C\cap N=1$, and $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation.
+    $C=\langle c_{1},\ldots,c_{m}\rangle\simeq A$, $G=C\cdot N$, $C\cap N=1$, and NN is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on NN by conjugation.
   3. 3)
@@
 Then $\phi(A)=C$, so $C\simeq A$.
-By Lemma 2.7, item 2), $N$ is the centralizer of the commutant $G^{\prime}$ of $G$, hence $\phi(N)=N$.
+By Lemma 2.7, item 2), NN is the centralizer of the commutant $G^{\prime}$ of $G$, hence $\phi(N)=N$.
 Clearly, $G=C\cdot N$ and $C\cap N=1$.
-The action of ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$ on $N$ comes from the conjugation of the elements $a_{i}$ on $N$, so $\phi$ maps this action to the action of ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ on $N$, and $N$ becomes a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on $N$ by conjugation.
+The action of ${\mathbb{Z}}[a_{1},a_{1}^{-1},\ldots,a_{m},a_{m}^{-1}]$ on NN comes from the conjugation of the elements $a_{i}$ on NN, so $\phi$ maps this action to the action of ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ on NN, and NN becomes a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$, when each $c_{i}$ acts on NN by conjugation.
 Hence, 1) implies 2).
@@
 Put $\Delta^{0}=1$ and let $\Delta^{i}$ be the $i$th-power of $\Delta$ for $i>0$, i.e., it is an ideal in ${\mathbb{Z}}A$ generated by all products of the type $(y_{1}-1)\ldots(y_{i}-1)$, where $y_{j}\in\{a_{1},\ldots,a_{m}\}$.
-Denote by $N^{\Delta^{i}}$ the ${\mathbb{Z}}A$-submodule of $N$ generated by all elements $u^{P}$, where $u\in N$ and $P\in\Delta^{i}$.
+Denote by $N^{\Delta^{i}}$ the ${\mathbb{Z}}A$-submodule of NN generated by all elements $u^{P}$, where $u\in N$ and $P\in\Delta^{i}$.
 Put $R={\mathbb{Z}}A$, $R_{i}={\mathbb{Z}}A/\Delta^{i}$, and $N_{i}=N/N^{\Delta^{i}}$.
@@
 To prove 1) recall that $G_{i}$ is generated as a normal subgroup by all simple left-normed commutators of the form $x_{1},\ldots,x_{i}]$ where $x_{k}\in\{a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}\}$, $k=1,\ldots,i$ (this is true for any group $G$ and $x_{i}$ from any fixed generating set of $G$).
-Since $N$ is abelian and $G^{\prime}\leq N$ any commutator $x_{1},\ldots,x_{i}]$ where $x_{s}\in N$ for some $s>2$ is equal to 1.
+Since NN is abelian and $G^{\prime}\leq N$ any commutator $x_{1},\ldots,x_{i}]$ where $x_{s}\in N$ for some $s>2$ is equal to 1.
 Therefore, $G_{i}$ is generated as a normal subgroup in $G$ by the commutators $b_{k},a_{j_{1}},\ldots,a_{j_{i-1}}]$, as claimed.
@@
 Proof.
-For any non-trivial $g\in N$, every element in $N$ commutes with $g$ and every element in $A$ does not commute with $g$.
-Since elery element in $G$ is a product of an element in $N$ and an element in $A$, this gives 1) and 2).
-Since $A$ is abelian, for any $a\in A$, $A\subseteq C_{G}(a)$.
+For any non-trivial $g\in N$, every element in NN commutes with $g$ and every element in AA does not commute with $g$.
+Since elery element in $G$ is a product of an element in NN and an element in AA, this gives 1) and 2).
+Since AA is abelian, for any $a\in A$, $A\subseteq C_{G}(a)$.
 And $a$ does not commute with any non-trivial $g\in N$, this gives 3).
@@
   • 1)
-    The subgroup $N$ is 0-definable in $G$ by the formula
+    The subgroup NN is 0-definable in $G$ by the formula
     $\phi_{N}(x)=\forall y\forall z[x,[y,z]]=1.$
   • 2)
-    The subgroup $A$ is definable in $G$ with parameters by the formula
+    The subgroup AA is definable in $G$ with parameters by the formula
     $\phi_{A}(x)=[x,a]=1,$
-    where $1\neq a$ is an arbitrary non-trivial element of $A$.
+    where $1\neq a$ is an arbitrary non-trivial element of AA.
 Proof.
@@
 for some $g\in A$, $g\neq 1$ and $z\in N$. We need to show that $g=a^{\gamma}$ for some $\gamma\in{\mathbb{Z}}$.
-Note first that in a free abelian group $A$ there is an automorphism $\theta\in\mathrm{Aut}A$ that maps $a$ to $a_{1}^{\beta}$ for some positive $\beta\in{\mathbb{N}}$.
-Replacing the basis $\{a_{1},\ldots,a_{m}\}$ of $A$ with basis $\{\theta(a_{1}),\dots,\theta(a_{m})\}$, if necessary, we may assume that $a=a_{1}^{\beta}$.
+Note first that in a free abelian group AA there is an automorphism $\theta\in\mathrm{Aut}A$ that maps $a$ to $a_{1}^{\beta}$ for some positive $\beta\in{\mathbb{N}}$.
+Replacing the basis $\{a_{1},\ldots,a_{m}\}$ of AA with basis $\{\theta(a_{1}),\dots,\theta(a_{m})\}$, if necessary, we may assume that $a=a_{1}^{\beta}$.
 If $z=1$ then $g=1=a^{0}$. Suppose $z\neq 1$ and write $z$ as $z=\prod_{i}b_{i}^{Q_{i}}$ for some Laurent polynomials $Q_{i}\in{\mathbb{Z}}A$. Then the equality (8) above takes the form
 $b_{1}^{g-1}=\Pi_{i}b_{i}^{Q_{i}(a_{1}^{\beta}-1)}.$
-Since $b_{1},\ldots,b_{n}$ is a basis of the free module $N$, one gets $Q_{i}=0$ for all $i>1$. Hence
+Since $b_{1},\ldots,b_{n}$ is a basis of the free module NN, one gets $Q_{i}=0$ for all $i>1$. Hence
 $g-1=Q_{1}(a_{1}^{\beta}-1)$
@@
 In particular, $g_{2}$ divides $g_{1}$.
-Since $g_{1}$ and $g_{2}$ do not have common divisors besides $1$ and 1-1, we deduce that $g_{2}=\pm 1$, and from $g=\frac{g_{1}}{g_{2}}$ we have $g_{2}=1$.
+Since $g_{1}$ and $g_{2}$ do not have common divisors besides 11 and 1-1, we deduce that $g_{2}=\pm 1$, and from $g=\frac{g_{1}}{g_{2}}$ we have $g_{2}=1$.
 It follows that $P_{1}=a_{1}^{\beta(k-1)}+\ldots a_{1}^{\beta}+1$ and $g=g_{1}=a^{k\beta}$ for some $k>0$.
@@
 We summarize this observation in the following corollary.
-Corollary 3.3.
+Corollary $3.$
 The following holds in $G$:
@@
 Replacing in this formula every occurence of the element $a_{i}$ by a variable $x_{i}$ we obtain a formula of group language that we denote by $Base_{top}(x_{1},\ldots,x_{m})$.
-Observe, that $G\models Base_{top}(a_{1},\ldots,a_{m})$, in fact, this it true for every basis $(a_{1}^{\prime},\ldots,a_{m}^{\prime})$ of $A$, because from the beginning $(a_{1},\ldots,a_{m})$ was chosen as an arbitrary basis of $A$.
+Observe, that $G\models Base_{top}(a_{1},\ldots,a_{m})$, in fact, this it true for every basis $(a_{1}^{\prime},\ldots,a_{m}^{\prime})$ of AA, because from the beginning $(a_{1},\ldots,a_{m})$ was chosen as an arbitrary basis of AA.
 Suppose that $G\models Base_{top}(c_{1},\ldots,c_{m})$ for some elements $c_{1},\ldots,c_{m}\in G$.
@@
 Denote by $d_{i}u_{i}$, where $d_{i}\in A,u_{i}\in N$, a generator of $C_{i}$.
-It is easy to see that $d_{1},\dots,d_{m}$ form a basis of $A$.
+It is easy to see that $d_{1},\dots,d_{m}$ form a basis of AA.
 Therefore, there is an automorphism $\psi\in Aut(A)$ such that $\psi(a_{i})=d_{i},i=1,\ldots,m$.
@@
 Clearly, $\psi$ extends linearly to an automorphism $\lambda_{\psi}$ of ${\mathbb{Z}}A$.
-Now, $\lambda_{\psi}$ gives rise to a semilinear bijection $\tilde{\psi}$ of the ${\mathbb{Z}}A$-module $N$, where for $u=\Pi b_{i}^{P_{i}}\in N,P_{i}\in{\mathbb{Z}}A$,
+Now, $\lambda_{\psi}$ gives rise to a semilinear bijection $\tilde{\psi}$ of the ${\mathbb{Z}}A$-module NN, where for $u=\Pi b_{i}^{P_{i}}\in N,P_{i}\in{\mathbb{Z}}A$,
 ${\tilde{\psi}}(u)=\Pi b_{i}^{\lambda_{\psi}(P_{i})}.$
@@
 Checking that it is a homomorphism of $G$ is straightforward if using the formula (3) of the multiplication in $G$.
-The isomorphism ${\tilde{\psi}}^{-1}$ maps $d_{i}u_{i}$ to $a_{i}v_{i}$, where $v_{i}={\tilde{\psi}}^{-1}(u_{i})\in N$, so $(a_{1}v_{1},\ldots,a_{m}v_{m})$ forms a basis of a free abelian group $A^{\prime}$, which is isopmorphic to $A$ via a map $\theta:a_{i}\to a_{i}v_{i},i=1,\ldots,n$.
-Note, that the elements $a_{i}v_{i}$ act by conjugation on $N$ precisely the same way as elements $a_{i}$, hence the map $au\to\theta(a)u$, for $a\in A,u\in N$ is an isomorphism of $G$.
+The isomorphism ${\tilde{\psi}}^{-1}$ maps $d_{i}u_{i}$ to $a_{i}v_{i}$, where $v_{i}={\tilde{\psi}}^{-1}(u_{i})\in N$, so $(a_{1}v_{1},\ldots,a_{m}v_{m})$ forms a basis of a free abelian group $A^{\prime}$, which is isopmorphic to AA via a map $\theta:a_{i}\to a_{i}v_{i},i=1,\ldots,n$.
+Note, that the elements $a_{i}v_{i}$ act by conjugation on NN precisely the same way as elements $a_{i}$, hence the map $au\to\theta(a)u$, for $a\in A,u\in N$ is an isomorphism of $G$.
 This shows that the tuples $(a_{1},\ldots,a_{m})$, $(a_{1},v_{1},\ldots,a_{m}v_{m})$, and $(d_{1}u_{1},\ldots,d_{m}u_{m})$ are automorphically equivalent in $G$.
 Recall that $c_{i}\in C_{i}=\langle d_{i}u_{i}\rangle$, so $c_{i}=(d_{i}u_{i})^{\gamma_{i}}$, for some $\gamma_{i}\in{\mathbb{Z}}$.
-Now we apply Lemma 3.2 where the basis $a_{1},\ldots,a_{m}$, hence the group $A$, is replaced with the basis $(d_{1}u_{1},\ldots,d_{m}u_{m})$ and the group $A^{\prime}$, and the element $a$ is replaced by $c_{i}\in A^{\prime}$.
+Now we apply Lemma 3.2 where the basis $a_{1},\ldots,a_{m}$, hence the group AA, is replaced with the basis $(d_{1}u_{1},\ldots,d_{m}u_{m})$ and the group $A^{\prime}$, and the element $a$ is replaced by $c_{i}\in A^{\prime}$.
 This gives us that the formula $Cyc(c_{i},y)$ defines in $G$ the cyclic subgroup $\langle c_{i}\rangle$ which is equal to $C_{i}$.
@@
 where $\bar{d}=(d_{1},\ldots,d_{m}),{\bar{a}}^{\bar{\gamma}}=(a^{\gamma_{1}},\ldots,a^{\gamma_{m}})$.
-In particular, the formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})=Prod_{A,2}(x_{1},x_{2},\bar{y},\bar{z})$ defines in $G$ the $\mathbb{Z}$}-exponentiation on $A$, that is for any $a,d\neq 1\in A$ and any $\gamma\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
+In particular, the formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})=Prod_{A,2}(x_{1},x_{2},\bar{y},\bar{z})$ defines in $G$ the $\mathbb{Z}$}-exponentiation on AA, that is for any $a,d\neq 1\in A$ and any $\gamma\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
 $G\models Prod_{A}(g,a,d,a^{\gamma})\Longleftrightarrow g=d^{\gamma}.$
@@
 $(au)^{k}=a^{k}u^{k}[v,a].$ (12)
-Moreover, $w=u^{k}$ is the only element in $N$ such that for every $a\in A$ the equation
+Moreover, $w=u^{k}$ is the only element in NN such that for every $a\in A$ the equation
 $(au)^{k}=a^{k}w[v,a]$ (13)
@@
 $w^{-1}u^{k}=[v_{1},a][v^{-1},a]=[v_{1}v^{-1},a].$
-Since $N$ is a free ${\mathbb{Z}}A$-module, one gets $w^{-1}u^{k}=1.$ Indeed, let $w^{-1}u^{k}=\Pi_{i}b_{i}^{P_{i}}$ and $v_{1}v^{-1}=\Pi_{i}b_{i}^{Q_{i}}$ for some $P_{i},Q_{i}\in{\mathbb{Z}}A$.
+Since NN is a free ${\mathbb{Z}}A$-module, one gets $w^{-1}u^{k}=1.$ Indeed, let $w^{-1}u^{k}=\Pi_{i}b_{i}^{P_{i}}$ and $v_{1}v^{-1}=\Pi_{i}b_{i}^{Q_{i}}$ for some $P_{i},Q_{i}\in{\mathbb{Z}}A$.
 Then for every $i$ $P_{i}=Q_{i}(a-1)$.
@@
 ∎
-Corollary 4.4.
+Corollary $4.$
 There is a formula $Exp_{N}(x_{1},x_{2},x_{3},y)$ of group theory language such that for any $a\neq 1\in A$, any $u\neq 1\in N$, and any $k\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
@@
 It follows from Corollaries 4, 6, and 7. ∎
-4.2 Interpretation of the ${\mathbb{Z}}A$-action on $N$ in $G$
-In this section, we show that the standard action of ${\mathbb{Z}}A$ on $N$ is interpretable in $G$ with parameters $a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}$.
+4.2 Interpretation of the ${\mathbb{Z}}A$-action on NN in $G$
+In this section, we show that the standard action of ${\mathbb{Z}}A$ on NN is interpretable in $G$ with parameters $a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}$.
 We fix an element $a=a_{1}\in A$ and use the interpretation ${\mathbb{Z}}_{a}$ of the ring of integers $\mathbb{Z}$} in $G$.
@@
 By induction on $m$. ∎
-Denote by $(N)^{I_{\bar{\alpha}}}$ the submodule of the ${\mathbb{Z}}A$-module $N$ generated by $N^{I_{\bar{\alpha}}}$.
+Denote by $(N)^{I_{\bar{\alpha}}}$ the submodule of the ${\mathbb{Z}}A$-module NN generated by $N^{I_{\bar{\alpha}}}$.
 Observe, that every element $u\in N^{I_{\bar{\alpha}}}$, as well as in $(N)^{I_{\bar{\alpha}}}$ admits a decomposition of the type $u=\Pi_{i}b_{i}^{U_{i}}$, where $U_{i}\in I_{\bar{\alpha}}$.
@@
 Lemma 4.7.
-There is a formula $\phi(x,\bar{y},\bar{z})$ of the group language such that for any basis $(a_{1},\ldots,a_{m})$ of $A$ and any tuple $(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the formula
+There is a formula $\phi(x,\bar{y},\bar{z})$ of the group language such that for any basis $(a_{1},\ldots,a_{m})$ of AA and any tuple $(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the formula
 $\phi(x,a_{1},\ldots,a_{m},a^{\alpha_{1}},\ldots,a^{\alpha_{m}})$
@@
 $\Pi_{i}b_{i}^{R_{i}P}=\Pi_{i}b_{i}^{S_{i}+U_{i}}$ (17)
-Since $N$ is free with basis $b_{1},\ldots,b_{n}$ it follows from (17) that for $i=1,\ldots,n$
+Since NN is free with basis $b_{1},\ldots,b_{n}$ it follows from (17) that for $i=1,\ldots,n$
 $R_{i}P=S_{i}+U_{i}$ (18)
@@
 for every $i=1,\ldots,n$. Hence $g^{P}=h$, as claimed. ∎
-Now we can prove that in the group $G$ the action of ${\mathbb{Z}}A$ on $N$ is definable.
+Now we can prove that in the group $G$ the action of ${\mathbb{Z}}A$ on NN is definable.
 Again, Laurent polynomials $Q\in{\mathbb{Z}}A$ are not elements of $G$, so we first represent them by their codes $k=\nu^{-1}(Q)$, and then represent the number $k$ as $a^{k}$ in the interpretation ${\mathbb{Z}}_{a}$ of $\mathbb{Z}$} in $G$.
@@
 $g^{P}=h\Longleftrightarrow G\models Act_{{\mathbb{Z}}[\bar{a}]}(g,h,a^{k},a).$
-This defines the action of the ring of ordinary polynomials ${\mathbb{Z}}[\bar{a}]$ on $N$.
-The last step is to get a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ that defines the action of the ring of the Laurent polynomials ${\mathbb{Z}}A$ on $N$.
+This defines the action of the ring of ordinary polynomials ${\mathbb{Z}}[\bar{a}]$ on NN.
+The last step is to get a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ that defines the action of the ring of the Laurent polynomials ${\mathbb{Z}}A$ on NN.
 Let $Q\in{\mathbb{Z}}A$.
@@
 Corollary 4.10.
-There is a formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ of group language such that for any $u\in N$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
+There is a formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ of group language such that for any $u\in N$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of NN, any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
 $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}\Longleftrightarrow G\models\Lambda(u,c_{1}^{\nu(\bar{P})},\bar{c},\bar{u}),$ (20)
@@
 Corollary 4.11.
-There is a formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ of group language such that for any $g\in G$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{\gamma}=(\gamma_{1},\ldots,\gamma_{m})\in{\mathbb{Z}}^{m}$, and any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
+There is a formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ of group language such that for any $g\in G$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of NN, any tuple $\bar{\gamma}=(\gamma_{1},\ldots,\gamma_{m})\in{\mathbb{Z}}^{m}$, and any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
 $g=c_{1}^{\gamma_{1}}\ldots c_{m}^{\gamma_{m}}u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}\Longleftrightarrow G\models\Lambda_{G}(g,c_{1}^{\bar{\gamma}},c_{1}^{\nu(\bar{P})},\bar{c},\bar{u}),$ (21)
@@
 By Corollary 4.10 for $u\in N$ and $P_{1},\ldots,P_{n}\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ the condition $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ from (21) is described by the formula $\Lambda_{N}$.
-Using this one can easily write a formula that states that the element $u$ such that $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ is unique in $N$, which means that $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
+Using this one can easily write a formula that states that the element $u$ such that $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ is unique in NN, which means that NN is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
 This proves the theorem.
 5 Bi-interpretability of $G$ and $\mathbb{Z}$}
-Fix a tuple of parameters $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively.
+Fix a tuple of parameters $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of AA and $B$ respectively.
 In this section we prove that $G$ and $\mathbb{Z}$} are bi-interpretable with parameters $\bar{p}$ in $G$.
@@
   1. 1)
-    Let $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively.
+    Let $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of AA and $B$ respectively.
 Then the group $G$ and the ring $\mathbb{Z}$} are bi-interpretable with each other through the interpretations $G\simeq\Delta({\mathbb{Z}})$ and ${\mathbb{Z}}\simeq\Gamma(G,\bar{p})$.
@@
 We begin by recalling some definitions from [13].
-For an abstract set $I$, we denote by $2^{I}$ the space of all subsets of $I$ endowed with the product topology (equivalently, the topology of pointwise convergence of the indicator functions).
+For an abstract set II, we denote by $2^{I}$ the space of all subsets of II endowed with the product topology (equivalently, the topology of pointwise convergence of the indicator functions).
 If $G$ is a group, we denote by $\mathcal{N}(G)$ the set of all normal subgroups of $G$ endowed with the subspace topology induced by the inclusion $\mathcal{N}(G)\subseteq 2^{G}$.
@@
   1. (a)
-    The center $Z(G)$ is a free group in the variety of abelian groups of exponent $2$, and $\{c_{k}\mid k\in\mathbb{Z}\}$ is a basis in $Z(G)$.
+    The center $Z(G)$ is a free group in the variety of abelian groups of exponent 22, and $\{c_{k}\mid k\in\mathbb{Z}\}$ is a basis in $Z(G)$.
   2. (b)
@@
 Proof.
-Let $N$ be the free nilpotent group of rank $2$ with basis $\{\tilde{a}_{i},\tilde{b}_{i}\mid i\in\mathbb{Z}\}$.
+Let NN be the free nilpotent group of rank 22 with basis $\{\tilde{a}_{i},\tilde{b}_{i}\mid i\in\mathbb{Z}\}$.
 It is well-known that $[N,N]=Z(N)$ is a free abelian group freely generated by the basic commutators $\tilde{a}_{i},\tilde{a}_{j}]$, $\tilde{b}_{i},\tilde{b}_{j}]$ for $j<j$, $i,j\in\mathbb{Z}$, and $\tilde{a}_{i},\tilde{b}_{j}]$ for $i,j\in\mathbb{Z}$ (see [29, Ch.
@@
 for all $i,j,\ell\in\mathbb{Z}$.
-It is straightforward to verify that $[N,N]/M$ is a free group in the variety of abelian groups of exponent $2$, freely generated by elements $c_{k}=[\tilde{a}_{0},\tilde{b}_{k}]M$, $k\in\mathbb{Z}$.
-Since $M$ is central in $N$, we can consider the quotient group $N/M$.
+It is straightforward to verify that $[N,N]/M$ is a free group in the variety of abelian groups of exponent 22, freely generated by elements $c_{k}=[\tilde{a}_{0},\tilde{b}_{k}]M$, $k\in\mathbb{Z}$.
+Since $M$ is central in NN, we can consider the quotient group $N/M$.
 We denote the images of $\tilde{a}_{i}$ and $\tilde{b}_{i}$ ($i,j\in\mathbb{Z}$) in $N/M$ by $a_{i}$ and $b_{j}$, respectively.
@@
 […]
-where $A$ is given by (23) and the indices $i$, $j$ independently range in $\mathbb{Z}$.
+where AA is given by (23) and the indices $i$, $j$ independently range in $\mathbb{Z}$.
 The map $\tilde{a}_{i}\mapsto\tilde{a}_{i+1}$, $\tilde{b}_{i}\mapsto\tilde{b}_{i+1}$ ($i\in\mathbb{Z}$) extends to an automorphism $N\to N$, which fixes $M$ setwise.
@@
 For instance, we have $G_{\emptyset}\cong\mathbb{Z}^{2}\,{\rm wr}\,\mathbb{Z}$.
-Lemma 6.6.
+Lemma $6.$
 For every non-empty subset $S\subseteq\mathbb{Z}$, the group $G_{S}$ is a central extension of the form
@@
 By Lemma 6.7, the relation $a,b_{0}^{t^{i}}]=1$ holds in $G_{T}$.
-On the other hand, the element $a,b_{0}^{t^{i}}]=c_{i}$ generates $Z(G_{S})\cong\mathbb{Z}_{2}$ by Lemma 6.6.
-In particular, $a,b_{0}^{t^{i}}]\neq 1$ in $G_{S}$.
+On the other hand, the element $a,b_{0}^{t^{i}}]=c_{i}$ generates $Z(G_{S})\cong\mathbb{Z}_{2}$ by Lemma $6.$ In particular, $a,b_{0}^{t^{i}}]\neq 1$ in $G_{S}$.
 By the definition of marked groups, this implies that $(G_{S},\{a,b_{0},t\})$ and $(G_{T},\{a,b_{0},t\})$ represent distinct elements of $\mathcal{G}$.
@@
 Let $\mathcal{C}=f(2^{\mathbb{Z}})$, where $f$ is the map from Lemma 6.9.
-Note that $\mathcal{C}$ consists of groups of the form (1) by Lemma 6.6.
-Since $2^{\mathbb{Z}}$ is compact and $f$ is continuous, $\mathcal{C}$ is compact; in particular, $\mathcal{C}$ is a closed subset of $\mathcal{G}$.
+Note that $\mathcal{C}$ consists of groups of the form (1) by Lemma $6.$ Since $2^{\mathbb{Z}}$ is compact and $f$ is continuous, $\mathcal{C}$ is compact; in particular, $\mathcal{C}$ is a closed subset of $\mathcal{G}$.
 Note also that $\mathcal{C}$ has no isolated points being the image of a space without isolated points under a continuous and injective map.